### v1

# AdaSemiCD: An Adaptive Semi-Supervised Change Detection Method Based on Pseudo-Label Evaluation

Abstract
Change Detection (CD) is an essential field in remote sensing, with a primary focus on identifying areas of change in bi-temporal image pairs captured at varying intervals of the same region by a satellite. The data annotation process for the CD task is both time-consuming and labor-intensive. To make better use of the scarce labeled data and abundant unlabeled data, we present an adaptive dynamic semi-supervised learning method, AdaSemiCD, to improve the use of pseudo-labels and optimize the training process. Initially, due to the extreme class imbalance inherent in CD, the model is more inclined to focus on the background class, and it is easy to confuse the boundary of the target object. Considering these two points, we develop a measurable evaluation metric for pseudo-labels that enhances the representation of information entropy by class rebalancing and amplification of confusing areas to give a larger weight to prospects change objects. Subsequently, to enhance the reliability of sample-wise pseudo-labels, we introduce the AdaFusion module, which is capable of dynamically identifying the most uncertain region and substituting it with more trustworthy content. Lastly, to ensure better training stability, we introduce the AdaEMA module, which updates the teacher model using only batches of trusted samples. Experimental results from LEVIR-CD, WHU-CD, and CDD datasets validate the efficacy and universality of our proposed adaptive training framework.
Index Terms:
Change Detection, Semi-supervised Learning, Remote Sensing, Mean-Teacher
publicationid: pubid: $0$–$00$ 2021 IEEE
I Introduction
Change detection (CD) has emerged as a significant research focus within the field of remote sensing in recent years. Its objective is to identify regions of interest that have experienced alterations in bi-temporal image pairs captured at varying times of the same geographical area using satellite technology. This method plays a crucial role in remote sensing data analysis and is of particular importance in various civilian sectors such as urban planning [1, 2], rural land management [3, 4], and disaster assessment [5, 6].
Figure 1: Comparison between SOTA semi-supervised methods and our AdaSemiCD on the CDD Dataset.
Given that the process of accurately annotating masks for change detection tasks is notably labor intensive, direct application of traditional supervised learning approaches, such as CNN [7, 3, 8, 9, 4] and Transformers [10, 11, 12], to a limited set of labeled data often results in limited performance. In response to these challenges, researchers have investigated a range of approaches such as Self-supervised learning(SSL)[13, 14],Unsupervised Change Detection (USCD) [15, 16, 17], Weakly Supervised Change Detection (WSCD) [18, 19, 20], and sample generation strategies [21, 22, 23, 24]. Although WSCD is cost-efficient, it is reliant on incomplete or inaccurate labels, which can introduce substantial errors and unpredictably noisy data. USCD, on the other hand, does not require labeled data and utilizes the intrinsic patterns present in the data, but it often faces challenges when tackling specific tasks like classification or detection. Sample generation strategies, which include data augmentation [23], generative adversarial networks (GANs) [22], and diffusion models [24], frequently necessitate the simulation or synthesis of additional data. However, when dealing with limited available samples, these methods may encounter constraints due to insufficient diversity in the generated data, which can diminish the model’s ability to generalize. As a result, semi-supervised change detection (SSCD) [25, 26, 27] emerges as a potentially more effective solution. The paradigm of semi-supervised learning [28, 29, 30] is to enhance CD performance by leveraging the limited available labels and the large volume of unlabeled samples. Usually, researchers generate pseudo-labels for the unlabeled data to act as guidance during training. These pseudo-labels are often temporary predictions with higher probabilities. The most prevalent approach is the Mean-Teacher [31] framework, which employ a teacher model to generate pseudo-labels that serve as guidance for the student model during the training process. The teacher model is subsequently updated using the exponential moving average (EMA) [32] of the student model. By training on a mix of limited data with actual labels and abundant data with pseudo-labels, the student model can learn more significant features, leading to noticeable enhancements in performance.
While these methods produce acceptable outcomes, significant problems persist: the model indiscriminately treats all samples, irrespective of their quality, and the training process lacks flexibility. Firstly, it’s evident that unlabeled samples may not always function as efficient ‘teachers’. Models frequently encounter difficulties in generating reliable high-quality pseudo-labels for intricate samples, which in turn introduces extra noise that can misguide the model’s training.
Subsequently, the EMA updating process does not take into account the quality of samples. Given that training batches can be biased or contain noise, dynamically determining the training update timing could contribute to the stability of the training process. These factors underscore the need for a more precise supervisory approach, failing which it could negatively impact the model’s training. In this study, we introduce an adaptive dynamic learning strategy, AdaSemiCD, designed to improve the accuracy of pseudo-labels and streamline the training process. Our framework incorporates the traditional semi-supervised training approach, augmented by two innovative functional modules, AdaFusion and AdaEMA. Initially, we utilize AdaFusion to suppress noise at the individual sample level, thereby enhancing the accuracy of pseudo-labels. Contrary to previous methods like Augseg [33] or CutMix [34] that relied on entirely random fusion regions, our AdaFusion technique proactively identifies the most uncertain region and substitutes it with reliable content from either labeled datasets or unlabeled datasets of superior quality. Following this, we dynamically adjust the rate of parameter updates in the teacher-student model via AdaEMA to ensure improved stability. Although the traditional EMA effectively mitigates fluctuations in model parameters, thereby boosting stability, it persists in uniformly updating after each training iteration, overlooking the model’s varying learning outcomes across different iterations when handling a range of training samples. If unlabeled samples contain an abundance of erroneous information, it can misdirect the model’s training. Therefore, our AdaEMA introduces an adaptive selection process for model-level parameter updates, enabling the model to fully integrate superior parameters.
In summary, the main contributions of this paper are as follows.
  • We propose an adaptive SSCD framework named AdaSemiCD, which dynamically improves the pseudo-labels as well as adjusts the training procedure with pseudo label quality assessment.
  • We propose an AdaFusion strategy to enhance unreliable unlabeled samples. The fusion region and the trusted contents are selectively chosen with the uncertainty map.
  • We propose an AdaEMA parameter update strategy, which updates the teacher model with a batch-wise pseudo-labels improving assessment.
  • Experimental results on publicly available datasets LEVIR-CD, WHU-CD, and CDD demonstrate the effectiveness of our method.
II Related Work
II-A Semi-supervised learning
The technique of semi-supervised learning (SSL) applies supervised learning to a small subset of labeled data and unsupervised learning to a larger pool of unlabeled data. SSL is mainly categorized into three strategies: consistent regularization (CR), self-training, and holistic methods. The last strategy combines the first two strategies within an SSL framework.
CR techniques are grounded in the concept of perturbed consistency, which utilizes the coherence between the model’s output after varying degrees of perturbed input data as a training constraint. The three consistent regularization frameworks, in ascending order, including: the ${\Pi}$-model [35], the Temporal-ensembling model [35], and the Mean-teacher (MT) model [32]. The ${\Pi}$-model’s double-branch network shares the weight; the Temporal-ensembling model amalgamates all the outputs in the time series, with each image’s pseudo-labels being the EMA of the previously generated results; the MT model carries out this smoothing operation at the model parameters level. This model has found application in subsequent semi-supervised research across various domains, such as Active-Teacher for semi-supervised object detection [36], [37, 38, 33, 39] for semi-supervised general semantic segmentation,  [40] for image classification, and  [41][42] for semi-supervised medical image segmentation. There has also been explored in the area of perturbation design, with [43] and [25] examining the image-level perturbation and feature-level perturbation of CR respectively.
In the realm of self-training methods, the authenticity of pseudo-labels is of paramount importance. This has led to extensive research into the effective selection of high-quality pseudo-labels for supervised learning. [44] employs a set probability threshold as a selection standard. ST++ [45] has developed a multi-tier self-training structure, where labels of high confidence are used for self-training repeatedly until all unlabeled samples have been utilized. [46] uses a constant entropy value as the filtering limit.
More often than not, CR and self-training are not used independently, but rather in conjunction, forming a comprehensive approach to semi-supervised learning. This is exemplified in [38, 33, 39, 40] and [45], as partially discussed in earlier sections.
II-B Semi-supervised Change Detection
Since annotating a large number of images for CD is time-consuming, recent methods mainly focus on the SSCD. In the realm of CR techniques, the incorporation of the mean teacher model in CD was first introduced by Bousias et al. [47]. However, the initial outcomes did not show considerable potential, as this SSCD method fell short when compared to a benchmark that exclusively used a restricted quantity of labeled data for entirely supervised learning. Despite increasing labeled data, this disparity continues to expand. Using this as a basis, Mao et al.[48] implemented minor and major improvements to the inputs of the teacher and student models, respectively. Furthermore, they formulated an extra teacher-virtual adversararial training component to further reduce the harmful effects of the pseudo label noise.
Additionally, other semi-supervised methods employ either a single model or a two-branch model with shared weights. Such as Sun et al. [49] introduce a siamese network. They incorporated additional self-training based on pseudo-labels, employing threshold filtering to eliminate low-quality pseudo-labels. The rationale behind this filtering lies in the potential noise introduced by pseudo-labels with low confidence, which could adversely affect SSL training. Hafner et al. [50] propose a dual-task SSCD framework that combines building segmentation and change detection, two closely related downstream tasks. They devised a novel consistency constraint between the two change detection masks produced by the Siamese segmentation network and the change detection network. Bandara et al. [25] explored feature-based perturbations of the regularization term, applying various data perturbations at the feature level to expand the distribution space of consistency constraints. This approach fully leverages the information embedded in unlabeled samples, and in recent work, Zhang et al. [27] imposed two constraints of class consistency and feature consistency on unlabeled datasets. By aligning the feature representations of unlabeled samples on varying and invariant classes, the model could learn from a feature space that is closer to the real distribution, this contributed to the renewal of the best performance record at that time.
Other methods predominantly utilize Generative Adversarial Network (GAN), which were initially proposed by Ian J. Goodfellow in 2014[51]. Some methods use GAN to learn the feature distribution space close to the real labeled data[26, 52, 53, 54]; The other part uses GAN to generate data samples[55, 56]; The more noteworthy research in the recent work is [18] proposed a new paradigm for change detection, unifying unsupervised, weakly supervised, regionally supervised, and fully supervised change detection into an end-to-end framework. While these endevavors have reported great success in SSCD, the highly unstable training of GAN makes hyperparameter adjustment become s challenging. Additionally, the issue of gradient disapperance frequently arises during the training phase. Morever, the strong discriminatory capaality of the discriminator can lead to imbalanced performance between the generator and discriminator of the GAN if no additional training techniques are implementrd. Consquently, achieving the idealized optimal scenario is challenging, making the pritical application of this method somewhat intricate. So we continue the semi-supervised framework of consistent and self-teinging.
III METHODOLOGY
Fig. 2 provides a comprehensive summary of our AdaSemiCD framework, aiming to enhance the SSCD performance by utilizing the scarce labels and vast quantity of unlabeled samples. We commence with a broad introduction of the framework, followed by an in-depth explanation of the uncertainty map used to assess our pseudo-labels, and finally present the specifics of AdaFusion and AdaEMA.
Figure 2: The overview of the proposed AdaSemiCD framework, and details of two novel modules. Our approach is based on the common teacher-student training pipeline for SSCD, and we suggest utilizing the AdaFusion module to produce samples with increased reliability and the AdaEMA module to enhance the efficiency of EMA update times.
III-A Overview of the AdaSemiCD Framework
The task of semi-supervised change detection can be described as follows: Given a labeled dataset ${D}_{l}=\left\{\left\{\smash{x_{a,i}^{l},x_{b,i}^{l}}\right\},y_{i}^{l}\right\}_{i=1}^{m}$ and an unlabeled dataset ${D}_{u}=\left\{x_{a,j}^{u},x_{b,j}^{u}\right\}_{j=1}^{n}$, where $\left\{\left\{\smash{x_{a,i}^{l},x_{b,i}^{l}}\right\},y_{i}^{l}\right\}$ represents the i-th pair of labeled image pairs and their ground truth, and $\left\{\smash{x_{a,i}^{u},x_{b,i}^{u}}\right\}$ represents the i-th pair of unlabeled image pairs, subscripts a and b are used to identify the images before and after phases respectively, and the number of samples with and without labeled datasets is n and m, $n\gg m$, the model $M$ is expected to not only extract meaningful information from ${D}_{l}$ but also capture a wider range of features by utilizing numerous unlabeled training samples in ${D}_{u}$ to improve the model’s generalization capabilities.
Model architecture: In this study, we employ the widely-used teacher-student framework for SSCD tasks. The network is composed of two components, a student model $M_{stu}$ and a teacher model $M_{tea}$ that shares the same structure. The student model is trained to extract significant features from a small number of labeled samples and a large volume of unlabeled samples, with the aid of optimization via gradient descent methods. Conversely, the teacher model $M_{tea}$ generates pseudo-labels to guide the student in assimilating unlabeled data, and the teacher model is refreshed using EMA methods.
Usually, samples undergo either weak $A_{w}(\cdot)$ or strong $A_{s}(\cdot)$ augmentation processes prior to being input into the network to maintain high-quality generalization performance.
Objectives: The objective is to minimize the supervised loss $\mathcal{L}_{s}$ on $D_{l}$ while also ensuring consistency on disturbed $D_{u}$ with a minimal $\mathcal{L}_{u}$. During training, samples are fed into the network in randomly shuffled batches, $\mathcal{B}_{l}$ and $\mathcal{B}_{u}$.
The loss $\mathcal{L}_{s}$ for labeled samples within a batch $B_{l}$ is calculated as the cross entropy (CE) of ground truth $y_{i}^{l}$ and it’s prediction $p_{i}^{l}$:
$\mathcal{L}_{s}=\frac{1}{\left|\mathcal{B}_{l}\right|}\sum_{i=1}^{\left|\mathcal{B}_{l}\right|}\mathrm{CE}\left({p}_{i}^{l},y_{i}^{l}\right),$ (1)
where $|\mathcal{B}_{l}|$ represents the mini-batch’s size, ${p}_{i}^{l}$ represents the CD result of the $i$-th pair images.
The loss $\mathcal{L}_{u}$ for unlabeled samples within a batch $B_{u}$ is quite similar. Here we use the pseudo-labels from $M_{tea}$ as supervision for the prediction from $M_{stu}$. And $\mathcal{L}_{u}$ is calculated as:
$\mathcal{L}_{u}=\frac{1}{\left|\mathcal{B}_{u}\right|}\sum_{j=1}^{\left|\mathcal{B}_{u}\right|}\mathrm{CE}\left(p_{s,j}^{u},{p}_{w,j}^{u}\right),$ (2)
where $|\mathcal{B}_{u}|$ represents the size of unlabeled image mini-batch, ${p}_{w,j}^{u}=M_{tea}^{\theta^{\prime}}(w_{a,j}^{u},w_{b,j}^{u})$ represents the detection result of the teacher model on the change of the i-th pair of unlabeled images weakly augmented, and ${p}_{s,j}^{u}={M_{stu}^{\theta}(s_{a,j}^{u},s_{b,j}^{u})}$ represents the detection result of the teacher model on the change of the i-th pair of unlabeled images by further strongly augmented.
To sum up, the total loss for the training procedure of AdaSemiCD is:
$\mathcal{L}=\mathcal{L}_{s}+\lambda(\cdot)\mathcal{L}_{u}$ (3)
Previous studies[40, 41] have employed a constant value for $\lambda(\cdot)$. However, we argue that this approach is not suitable for single-stage semi-supervised learning algorithms. During the initial training phases, the pseudo-labels generated by our model for unlabeled samples are highly unreliable; excessive reliance on unsupervised training at this stage can introduce significant noise. Conversely, as training progresses into the middle and later stages, the model increasingly learns from the limited labeled data, leading to an improvement in the quality of the pseudo-labels. This justifies a reduction in the proportion of supervised training relative to unsupervised training. To mitigate overfitting and enhance the feature space, it is essential to implement a controlled process to dynamically adjust the contribution of these two components in the loss function. As the pseudo-labels in the starting stage are quite unstable, we use a ramp-up process to control the training pace of the unlabeled part. $\lambda(\cdot)$ denotes the time varing weight of unsupervised loss, which is a function changing with training step and used to dynamically adjust the proportion of supervised and unsupervised training in different stages.
$\lambda(\cdot)=w_{max}\times e^{-\phi\times\left(1-iter_{cur}/iter_{max}\right)^{2}}$ (4)
where $w_{max}$ and $\phi$ are hyperparameters, $w_{max}$ representing the maximum weight of unsupervised loss, $\phi$ is used to control the severity of the ramp-up; $iter_{cur}$ represents the current iteration cycles; $iter_{max}$ is the total number of ramp-up cycles, is obtained by multiplying $\gamma$ and the total number of training iterations, and $0\textless\gamma\textless 1.0$; After the ramp-up process, the weight of unsupervised losses remains $w_{max}$ and does not change. In the early stage of training, this weight is relatively low, and unsupervised training plays a negligible role, but in the middle and later stages of training, this weight gradually increases and the unsupervised loss after weighting exceeds the supervised loss, so unsupervised training accounts for the main part.
Training strategy: The student network’s parameters $\theta$ are adjusted by Stochastic Gradient Descent(SGD) technique to reduce the overall loss $\mathcal{L}$, while the teacher network’s parameters $\theta^{\prime}$ are updated by the exponential moving average of the student model’s parameters $\theta$ over the time series, as depicted in Eq. 5. The hyperparameter $\beta$ acts as a momentum parameter. The larger the value of $\beta$, the broader the moving metric average window becomes. Typically, $\beta$ is chosen to be a number near 1.0, for instance, 0.996 in this study.
$\theta^{\prime}=\beta\theta^{\prime}+(1-\beta)\theta$ (5)
Proposed modules: The essence of semi-supervised learning lies in the quality of pseudo-labels. Nevertheless, it’s apparent that the previously mentioned procedure does not take into account the diverse effects of individual samples on training. This paper concentrates on two elements that are directly related to the generation of pseudo-labels: the initial pair of unlabeled images, and the efficiency of the pseudo-label-generation network, also known as the teacher model, in identifying changes. To offer more dependable supervisory data to unlabeled information and reduce training uncertainty, we’ve designed an adaptive training strategy to tackle these two key issues. Our initial proposal is to develop a metric that can measure the uncertainty of pseudo-labels, serving as the basis for adaptive modifications. Following that, we suggest making adaptive changes at the image level to the unlabeled training samples and suitably integrating reliable contents. Furthermore, it includes applying adaptive and selective EMA updates to the teacher network during the training phase to limit variations, thus producing more consistent and superior quality pseudo-labels. These efforts are elaborated in the subsequent sections.
III-B Pseudo-label qualification metric
To enhance the effectiveness of pseudo-labels by accurately gauging their quality, a crucial step is the implementation of an evaluation metric. This metric will be instrumental in identifying reliable labels and determining the utility of each training sample. Unlike the scenario with labeled image pairs where real labels serve as a benchmark for assessing the model’s inference quality through average intersection over union, F1 scores, and so forth, pseudo-labels lack any such reference points and can only be compared to themselves. Therefore, we have devised a quantifiable evaluation metric specifically for pseudo-labels, designed to increase the total information entropy while taking into account factors like category imbalance and areas of confusion.
To evaluate the quality of pseudo-labels, a frequently employed technique for this is the computation of information entropy [57],
$E(x_{i})=-P\left(x_{i}\right)\log_{2}P\left(x_{i}\right),$ (6)
where $P\left(x_{i}\right)$ is the output probability of a trained model on sample $x_{i}$. Our conviction is that a lower information entropy in the predicted value corresponds to a greater trustworthiness of the prediction outcome. On the contrary, a higher information entropy leads to a more varied prediction outcome, a more evenly distributed prediction probability on the same pixel, and a diminished distinction in the category.
Figure 3: Class statistics of common CD datasets. The proportion of the changed/unchanged categories is extremely unbalanced.
In our change detection task, the direct application of entropy often does not yield good results due to the significant challenge posed by category imbalance. As illustrated in Fig. 3, ten publicly available change detection datasets were evaluated, namely HRCUS-CD[58], LEVIR-CD[59], LEVIR-CD+[59],WHU-CD[60], EGY-CD[61], GZ-CD[26], CL-CD[62], CDD[63], SYSU-CD[64] and DSIFN-CD[65]. The ratios of the changed and unchanged categories are highly skewed. This significant category imbalance can lead to our model inadequately learning the target categories during training while disproportionately capturing feature distributions from background categories. As a result, there is an increased tendency to classify pixels as belonging to the background class during inference. Consequently, although the overall accuracy (OA) of the model appears high, the performance on the Intersection over Union (IoU) across various categories remains unsatisfactory. To minimize the influence of this category imbalance when assessing the quality of pseudo-labels, we assigned different weights to the two categories when computing information entropy as Eq. 7.
$E^{\prime}\left(X\right)=w_{1}\times E\left(x_{i}\right)[0]+w_{0}\times E\left(x_{i}\right)[1],$ (7)
where $w_{0}$ and $w_{1}$ represent the proportions of pixels in the current batch that belong to the unchanged and changed categories, respectively. The formulas for their calculation are given by Eq. 8 and Eq. 9.
$w_{0}=\frac{\sum_{i=1}^{|Bu|}\sum_{k=1}^{H\times W}{(P(x_{i})==0)}}{{|Bu|}\times H\times W}$ (8)
$w_{1}=\frac{\sum_{i=1}^{|Bu|}\sum_{k=1}^{H\times W}{(P(x_{i})==1)}}{{|Bu|}\times H\times W}$ (9)
Moreover, pixels from diverse areas contribute in varying manners. Pixels of greater importance, like those in boundary, target, and regions similar to the background, who are offenly predicted with large uncertainty, should receive more emphasis. Initially, we enhance the impact of these essential regions by computing the absolute value of the disparity between the prediction probabilities of the two categories.
$D\left(x_{i}\right)=abs(P\left(x_{i}\right)[1]-P\left(x_{i}\right)[0])$ (10)
Here, $abs$ signifies absolute value, a measure taken to prevent the inconsistencies that may arise from varying changes both prior to and following a phase. Then doing pixel-level product operation with information entropy, we get the uncertainty map of image $(x_{i})$,
$U\left(x_{i}\right)=1-D\left(x_{i}\right)\cdot E^{\prime}\left(x_{i}\right)$ (11)
Clearly, in pixel locations with low information entropy, this process won’t result in notable alterations, whereas in pixel locations with high information entropy, the absolute value of the predictive probability post-subtraction will be significantly reduced, and the immediate amalgamation will diminish its importance. Consequently, we adopt the inverse of rectified information entropy as its depiction.
To sum up, with the goal of evaluating the quality of pseudo-labels for change detection, we put forward a measurable calculation metric $U$ to gauge uncertainty, which enhances the overall information entropy considering factors like category imbalance and focus on areas of confusion.
III-C AdaFusion: Adaptive sample fusion
Image fusion is frequently employed to augment samples and boost generalization, with CutMix[34] and MixUp[66] being typical examples. In this study, our aim is to leverage image fusion techniques to exclude the unreliable areas in a training sample. This procedure involves two steps: region selection, which decides the location of the operation, and image selection, which decides the content to be used.
Adaptive selection of fusion region. Contrary to the conventional CutMix technique, which arbitrarily chooses blending regions, our approach is more refined. Initially, we would select a bounding box with random size, then slide the window, and choose the area where the total of uncertainty is highest as the region to be combined, using Eq. 11.
These regions are frequently boundaries or intricate areas where it’s challenging to identify targets. The CutMix fusion can guarantee the variety of samples and simultaneously decrease the unsupervised noise.
Adaptive selection of fusion contents. For the region of maximum uncertainty, we have the option to choose a substitute from its corresponding samples from the label section $B_{l}$ or the unlabeled section with higher probabilities $B_{u}$. This strategy prevents overuse of only labeled samples and further mitigates the risk of overfitting. The focus of the content here is confined to the contrast between labeled and unlabeled image pairs, rather than a more detailed distinction (such as the level of the image or the level of the object). This is because we suggest that too much human intervention could potentially undermine the model’s capacity to generalize. The choice of fusion content is mainly accomplished by adaptively adjusting a threshold that determines whether to fuse with a labeled image pair. We directly use the computed uncertainty as the threshold, and if the randomly generated probability exceeds the total uncertainty of the sample, the labeled images in the training batch are randomly selected as the fusion content. If not, other unlabeled image pairs in the batch are randomly selected. It is easy to understand that the higher the sample’s reliability, the better the quality of the pseudo-label. We consider it reliable enough to avoid the necessity for fusion with the limited labeled images, which could potentially increase the likelihood of overfitting. Samples with higher uncertainty come with more noise, and the addition of new content can significantly reduce the noise density.
III-D AdaEMA: Adaptive EMA update strategy
Within the teacher-student framework, the teacher model is obtained by integrating the moving exponential average of the student model across the time series. This approach is more stable and reliable compared to dual models with shared weights. However, our expectation for Mean-Teacher is to reach the optimal state of coevolution, where the teacher model is a cumulative representation of the evolving student model. The crucial element in achieving this co-evolution is whether the student model progresses or regresses each time the teacher model gets updated. So, how can we assess the state of the student model? This question is inherently related to the validation phase of the training process. Typically, after training for several epochs, we evaluate the current model, feed samples from the validation set into the model for inference, and compare the outcomes with the actual labels to determine accuracy. However, if we perform such validation after every iteration, it becomes highly computationally intensive, leading to a significant increase in training time. Lowering the number of samples in the validation set (to just a few pairs) might address the feasibility issue, but it also presents new challenges when the sample size is too small to thoroughly assess the model’s performance. So, is there a middle ground that can merge these two concepts?
Algorithm 1 The AdaEMA algorithm.
0:    Student model $M_{stu}^{\theta}$, Teacher model $M_{tea}^{\theta^{\prime}}$ The set of training samples for current batch =$\mathcal{B}=\left\{\mathcal{B}_{l},\mathcal{B}_{u}\right\}$;
0:    Updated Teacher model, $M_{tea}^{\hat{\theta}}$;
1:  Calculate the loss $\mathcal{L}_{s}$ on labeled samples $\mathcal{B}_{l}$ with Eq. 1
2:  Calculate the loss $\mathcal{L}_{u}$ on unlabeled samples $\mathcal{B}_{u}$ with Eq. 2
3:  Update the student model $M_{stu}^{\theta}$ to $M_{stu}^{\theta^{\prime}}$ by SGD to minimize the total loss following Eq. 3
4:  Calculate the uncertainty ${U}_{tea}$ of pseudo-labels generated on $\mathcal{B}_{u}$ by the teacher model $M_{tea}^{\theta^{\prime}}$according to Eq.11
5:  Calculate the uncertainty ${U}_{stu}$ on $\mathcal{B}_{u}$ by updated student model $M_{stu}^{\theta^{\prime}}$ according to Eq.11
6:  Calculate the upper bound on the probability of update $\tau$ according to Eq. 12 and Eq. 13;
7:  if $random(\cdot)<\tau$ then
8:     Update the teacher model to $M_{tea}^{\hat{\theta}}$ by Eq. 5
9:  else
10:     $M_{tea}^{\hat{\theta}}=M_{tea}^{\theta^{\prime}}$
11:  end if
12:  return  $M_{tea}^{\hat{\theta}}$.
At each training phase, we initially update the student model’s parameters, denoted as $M_{stu}^{\theta}$, following the training strategy described in section III-A, resulting in $M_{stu}^{\theta^{\prime}}$. Next, we evaluate both the updated student model $M_{stu}^{\theta^{\prime}}$ and the teacher model $M_{tea}^{\theta^{\prime}}$ on the current batch of unlabeled training samples $\mathcal{B}_{u}$. We determine the ensuing uncertainty map using Eq. 11, and denote these as $U_{tea}$ and $U_{stu}$. The progression is marked by changes in uncertainty, as illustrated in Eq. 12.
$\varepsilon=\frac{\sum{U_{stu}}-\sum{U_{tea}}}{|B_{u}|}.$ (12)
If $\varepsilon$ is below zero, the model either degrades or oscillates. In contrast, when $\varepsilon$ is zero or higher, the model progresses. Ultimately, only the student models that have evolved successfully are selected to refresh the teacher model, with the maximum probability of update $\tau$ being derived as presented in Eq. 13.
$\tau=\left\{\begin{array}[]{cl}\frac{1}{iter^{2}+\epsilon}&,\quad\varepsilon\leq 0\\ 1.0&,\quad\varepsilon>0\\ \end{array}\right.,$ (13)
Here, $\epsilon=1e-5$ is used to prevent the divisor from being zero, and $iter$ denotes the present iteration count. If the model parameters are not selected for updating the Teacher model, they are not discarded outright. Instead, they are given a lower probability of being updated. With this upper bound regulation (ranging from 0 to 1), some randomness is introduced to determine the likelihood of updating the teacher network’s parameters. Further details are available in Algorithm 1.
IV Experiement
This section pertains to the experimental validation we performed. Initially, we present the assessment standards for experimental outcomes and the experimental datasets. Subsequently, we executed ablation experiments to validate the significance of each component in the methodology. Ultimately, we draw comparisons with the cutting-edge semi-supervised change detection techniques, using measurable experimental metrics and perceptible visual outcomes.
IV-A Experimental setup
IV-A1 Datasets
Our method is empirically tested on ten benchmark datasets, namely the LEVIR-CD [59], LEVIR-CD+ [59], WHU-CD[60], EGY-CD[61], HRCUS-CD[58], Change Detection Dataset(CDD)[63], GZ-CD[26], DSIFN-CD[65], SYSU-CD[64] and CL-CD[62]. Among them, these datasets cover different resolutions (0.03m-2.0m), different data sizes (2400-20000 pairs), different annotation categories (binary or multiclass), and different time spans between image pairs (1-16 years), as summarized in the Table I.
TABLE I: Ten publicly available change detection datasets used in the experiment.
Category Datesets Spatial Resolution Size Annotated Samples Time Spans Download
LEVIR-CD [59] 0.5m 1024 $\times$ 1024 637 5 to 14 years Link
LEVIR-CD+ [59] 0.5m 1024 $\times$ 1024 985 5 to 14 years Link
WHU-CD [60] 0.2m 15354$\times$32507 1 2012 to 2016 Link
GZ-CD [26] 0.55m Varying 19 2006 to 2019 Link
EGY-BCD [61] 0.25m 256 $\times$ 256 6091 2015 to 2022 Link
Building HRCUS-CD [58] 0.5m 256 $\times$ 256 11388 Varing Link
CDD[63] 0.03m-1.0m 256$\times$256 16000 Varing Link
DSIFN-CD [65] Unknown 512$\times$512 3940 Unknown Link
SYSU-CD[64] 0.5m 256$\times$256 20000 2007 to 2014 Link
Multiclass CL-CD [62] 0.5-2.0m 512$\times$512 600 2017 to 2019 Link
[…]
IV-A2 Evaluation Metrics
To facilitate a straightforward comparison with the latest cutting-edge techniques, we employed intersection over union (IoU) scores and overall accuracy (OA) on evaluation datasets. Given the severe imbalance in CD categories and our primary concentration on the changed area, we use the IoU of the change category $IoU^{c}$, the calculation formula are as follows:
$\mathrm{IoU}=\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FP}+\mathrm{FN}}$ (14)
$\mathrm{OA}=\frac{\mathrm{TP}+\mathrm{TN}}{\mathrm{TP}+\mathrm{FP}+\mathrm{TN}+\mathrm{FN}}$ (15)
Where $\mathrm{TP}$ represents the positive sample correctly predicted (the correct changing pixel), $\mathrm{TN}$ refers to the negative sample correctly predicted (the correct unchanged pixel), and $\mathrm{FP}$ denotes the positive sample wrongly predicted (the unchanged pixel wrongly detected), $\mathrm{FN}$ represents the negative sample wrongly predicted (the pixel missed as the unchanged pixel). For both metrics, the larger the value, the better the change detection performance of the model. In the change detection task, OA is high overall due to binary classification and unbalanced category envy.
IV-A3 Implementation Details
The common components of this experiment were tested under the same conditions as the methods we intend to compare them with. To prevent network structure from causing interference, we utilized ResNet50+PPM as our change detection network follow [25] and [27]. We set the initial learning rate at 0.01, which linearly decreases to 1e-4 with a momentum of 0.9, and trained all methods using the SGD optimizer. All models underwent 80 epochs of training, with each batch having a mini-batch size of 8 for both labeled and unlabeled training sets. Furthermore, all the enhancements employed were identical to those used in FPA[27], which include random flip, random resize ranging from 0.5 to 2.0, and random crop among the weak augmentations, along with the nine strong augmentations cited in [67]. Given the limited amount of labeled data, the model is highly sensitive to the pseudo-labels’ filtering threshold, leading us to set 0.95 as the threshold for all models in the experiments conducted on the ten datasets and we follow the experience in [25] to set $\phi$ to 5. All our experiments were carried out using PyTorch on four NVIDIA GeForce RTX 3090 GPUs.
IV-B Compare with the SOTA methods
In order to prove the superiority of our proposed method, we compare it with several of the most advanced semi-supervised change detection methods. [26], [25] and [27] refresh the last semi-supervised CD performance in chronological order. In addition, we also compare with two semi-supervised semantic segmentation modding methods[37] and [68].
IV-B1 Quantitative results
We conducted a complete experiment on ten datasets, Table II and Table III report our experimental results on the Building CD datasets and mutilclass CD datasets respectively. Sup.Only is the result of supervised training only on a limited subset of labeled dataset, while Oracle is the result of fully supervised training on the entire training dataset. It is worth noting that our method achieves the state-of-the-art performance in almost all partition settings across the all datasets. To be sure, all the semi-supervised CD methods performed better than the supervised methods in the corresponding partition setting in most cases, which proves the effectiveness of the semi-supervised methods in learning from a large number of unlabeled training samples, at the same time, the lead achieved by our method proves that our adaptive design for how to learn from unlabeled samples to more correct guidance is of practical significance.
[…]
In addition, from the experimental results, the overall accuracy(OA) of all methods on these binary change detection datasets is very high, because the background class accounts for the majority. OA is often directly corresponding to $IoU_{c}$, and a slight improvement in OA is reflected in a relatively significant improvement in $IoU_{c}$. Our AdaSemiCD improves OA in almost all datasets because it greatly eliminates noise and the guidance of false signals.
TABLE II: The average quantitative metrics of different CD methods on building change detection datasets. The highlighted parts in blue are the best results, and the underlined ones are the second best results.
Dataset Method 5% 10% 20% 40%
$IoU^{c}$ OA $IoU^{c}$ OA $IoU^{c}$ OA $IoU^{c}$ OA
LEVIR-CD Sup. only 61.0 97.60 66.8 98.13 72.3 98.44 74.9 98.60
AdvEnt[68] 66.1 98.08 72.3 98.45 74.6 98.58 75.0 98.60
s4GAN[37] 64.0 97.89 67.0 98.11 73.4 98.51 75.4 98.62
SemiCDNet[26] 67.6 98.17 71.5 98.42 74.3 98.58 75.5 98.63
RCR[25] 72.5 98.47 75.5 98.63 76.2 98.68 77.2 98.72
FPA[27] 73.7 98.57 76.6 98.72 77.4 98.75 77.0 98.74
AdaSemiCD 77.7 98.78 79.4 98.87 80.3 98.92 80.6 98.93
Oracle $IoU^{c}$=77.9 and OA=98.77
LEVIR-CD+ Sup. only 52.0 97.72 58.4 98.06 66.1 98.31 66.2 98.42
AdvEnt[68] 52.2 97.68 59.9 98.11 65.9 98.37 68.0 98.51
s4GAN[37] 46.5 97.25 51.4 97.66 62.8 98.18 67.2 98.46
SemiCDNet[26] 52.6 97.66 60.7 98.24 64.8 98.37 66.1 98.38
RCR[25] 64.9 98.25 67.5 98.45 68.5 98.52 68.4 98.51
FPA[27] 64.6 98.30 67.3 98.40 70.3 98.64 69.0 98.59
AdaSemiCD 66.7 98.49 68.8 98.51 70.6 98.63 70.9 98.64
Oracle $IoU^{c}$=70.5 and OA=98.63
WHU-CD Sup. only 50.0 97.48 55.7 97.53 65.4 98.20 76.1 98.94
AdvEnt[68] 55.1 97.90 61.6 98.11 73.8 98.80 76.6 98.94
s4GAN[37] 18.3 96.69 62.6 98.15 70.8 98.60 76.4 98.96
SemiCDNet[26] 51.7 97.71 62.0 98.16 66.7 98.28 75.9 98.93
RCR[25] 65.8 98.37 68.1 98.47 74.8 98.84 77.2 98.96
FPA[27] 66.3 98.45 57.4 97.69 62.5 98.48 73.1 98.69
AdaSemiCD 67.8 98.62 70.8 98.70 74.7 98.86 79.6 99.13
Oracle $IoU^{c}$=85.5 and OA=99.38
GZ-CD Sup. only 47.5 93.56 51.4 94.26 58.0 95.65 66.3 96.62
AdvEnt[68] 48.6 94.39 50.9 94.89 60.2 95.79 66.2 96.58
s4GAN[37] 50.8 94.38 52.4 94.98 60.8 95.94 64.2 96.39
SemiCDNet[26] 48.4 93.58 49.7 94.79 59.0 95.66 66.3 96.57
RCR[25] 50.8 93.82 50.8 94.69 62.5 96.07 67.8 96.61
FPA[27] 51.2 93.92 58.9 95.78 63.1 96.26 68.2 96.82
AdaSemiCD 51.6 94.56 57.1 95.57 62.4 96.21 68.0 96.75
Oracle $IoU^{c}$=69.0 and OA=96.93
EGY-CD Sup. only 49.8 95.73 54.6 96.38 61.4 96.83 65.1 97.25
AdvEnt[68] 52.7 96.01 57.8 96.58 62.6 96.86 64.0 97.19
s4GAN[37] 52.9 95.94 58.6 96.50 64.7 97.09 64.9 97.27
SemiCDNet[26] 52.4 96.00 57.9 96.31 62.8 96.95 63.8 97.19
RCR[25] 58.1 96.50 61.9 96.77 63.9 97.08 64.2 97.18
FPA[27] 57.5 96.52 60.1 96.86 65.2 97.25 65.7 97.34
AdaSemiCD 59.0 96.55 60.5 96.80 65.0 97.20 67.4 97.39
Oracle $IoU^{c}$=67.6 and OA=97.54
HRCUS-CD Sup. only 29.5 98.11 36.0 98.45 43.4 98.68 48.9 98.84
AdvEnt[68] 29.1 98.11 36.9 98.40 42.5 98.61 48.8 98.71
s4GAN[37] 25.0 97.86 28.2 98.24 40.1 98.62 50.3 98.85
SemiCDNet[26] 28.4 98.00 34.7 98.44 44.1 98.68 48.5 98.74
RCR[25] 36.1 98.36 42.1 98.69 45.3 98.76 49.6 98.66
FPA[27] 35.2 98.37 43.7 98.65 46.7 98.82 51.2 98.81
AdaSemiCD 37.8 98.59 42.6 98.70 48.1 98.84 50.8 98.87
Oracle $IoU^{c}$=59.0 and OA=99.06
TABLE III: The average quantitative metrics of different CD methods on multiclass change detection datasets. The highlighted parts in blue are the best results, and the underlined ones are the second best results.
Dataset Method 5% 10% 20% 40%
$IoU^{c}$ OA $IoU^{c}$ OA $IoU^{c}$ OA $IoU^{c}$ OA
CDD-CD Sup. only 60.4 94.25 67.9 95.46 75.6 96.59 82.3 97.56
AdvEnt[68] 63.3 94.65 71.2 96.01 79.3 97.14 82.9 97.66
s4GAN[37] 62.3 94.69 71.0 95.94 79.0 97.10 82.8 97.63
SemiCDNet[26] 63.5 94.68 71.2 95.99 79.1 97.13 82.8 97.63
RCR[25] 67.6 95.40 75.5 96.57 80.2 97.26 82.7 97.61
FPA[27] 68.9 95.66 74.9 96.55 79.7 97.20 81.1 97.37
AdaSemiCD 70.1 95.89 77.3 96.89 82.1 97.56 83.9 97.80
Oracle $IoU^{c}$=87.8 and OA=98.10
DSIFN-CD Sup. only 34.8 78.34 38.9 83.41 40.2 87.00 39.6 87.00
AdvEnt[68] 31.8 77.83 36.3 83.86 40.8 85.92 37.4 86.31
s4GAN[37] 36.6 84.10 34.8 86.87 37.9 87.69 40.1 86.52
SemiCDNet[26] 33.6 78.60 37.9 84.18 39.1 86.77 39.1 87.05
RCR[25] 26.7 83.78 32.9 86.05 40.8 86.70 36.7 86.08
FPA[27] 39.2 84.27 38.5 87.12 36.0 87.41 35.8 86.50
AdaSemiCD 36.9 80.46 39.2 82.94 41.1 85.45 45.1 87.12
Oracle $IoU^{c}$=37.1 and OA=86.82
SYSU-CD Sup. only 62.9 89.57 64.4 90.18 66.0 90.82 66.4 90.93
AdvEnt[68] 61.2 89.36 64.5 90.18 65.7 90.35 68.3 91.24
s4GAN[37] 64.4 90.02 66.5 90.48 66.9 90.26 68.2 91.51
SemiCDNet[26] 61.7 89.32 64.8 90.25 66.7 90.97 67.0 91.08
RCR[25] 62.5 89.76 66.0 90.75 64.1 90.22 65.3 90.56
FPA[27] 67.7 90.95 68.3 91.09 70.1 92.01 69.3 91.97
AdaSemiCD 67.5 91.16 68.7 91.59 70.1 92.03 69.9 91.90
Oracle $IoU^{c}$=77.9 and OA=98.77
CL-CD Sup. only 18.1 91.90 31.4 92.42 37.2 93.32 45.9 94.98
AdvEnt[68] 24.3 92.13 33.2 93.01 37.6 93.59 42.9 94.06
s4GAN[37] 22.1 92.00 26.6 93.09 37.4 93.59 43.4 93.87
SemiCDNet[26] 24.0 92.20 28.3 93.42 36.2 92.41 45.3 94.22
RCR[25] 27.1 91.63 32.8 92.99 36.4 93.07 48.5 94.94
FPA[27] 29.0 91.00 38.2 93.37 39.6 93.88 43.1 94.15
AdaSemiCD 30.6 92.52 33.5 92.40 41.6 94.21 49.1 95.85
Oracle $IoU^{c}$=77.9 and OA=98.77
Multiclass CD Datasets: As shown in TableIII, our AdaSemiCD achieved SOTA performance on almost all datasets and all settings, just the improvement was less than on the binary building change detection datasets. This is due to the complexity of the task itself, and one of the most obvious is that in some cases, some of the previous semi-supervised change detection methods failed and performed worse than the baseline under full supervision. However, our semi-supervised CD approach can still leverage additional unlabeled samples to improve performance.
Another point worth noting is that in these multiclass change detection datasets, the overall accuracy of all methods is reduced, and OA and $IoU_{c}$ are no longer directly corresponding. One of the phenomena is that $IoU_{c}$ reaches the highest, but OA does not, which is due to its different emphasis. OA pays more attention to the global accuracy, that is, the detection of background categories. $IoU_{c}$ is more concerned with the detection of the target class, which is more important in CD than the background class. As you can see in DSIFN-CD, our approach has achieved good results on the $IoU_{c}$.
IV-B2 Qualitative results
To provide a clearer illustration of the change detection capability of our proposed method, we conducted visualizations on three different datasets. Fig. 4 showcases an example of the visualizations on the WHU-CD test sets. It is apparent that on both datasets, our approach has notably mitigated the common issues of missed and false detections. In challenging scenarios, our method could still effectively identify the areas of change that were of interest to us. Regarding the image pairs in the lower part of Fig. 4, there is significant interference information present between the images. Many of the detected changes are unrelated to the buildings of interest. The absence of adequate supervision information makes it challenging to mitigate such interference, leading to decreased model performance. Additionally, the task is further complicated by the detection of small and densely changing areas, which proves to be difficult for the model.
Alternative semi-supervised models may have been trained to detect errors because of guidance issues in this complicated scenario. Consequently, it can be seen that in such instances the semi-supervised method performs even worse than the supervised method trained solely on labeled data. Our model, on the other hand, adaptively excludes these noises during the training phase and integrates parameters from the model that has completed its evolution through the training process. In these intricate cases, the quality of pseudo-labels improves progressively until achieving a reliable level, furnishing an accurate signal for model training. Therefore, it is evident that our model excels in these complex regions (highlighted by boxes in Fig. 4). This marks the foundation of our proposed method.
Figure 4: Visualizations of different models on the WHU-CD dataset. Pre and Post represent the image of the front and back temporal respectively, and GT is the Ground Truth.
IV-C Ablation study
In this section, we conduct some ablation studies, mainly to verify the role of each module in our proposed approach and to explore some of our hyperparameter selections in the experiment.
Effectiveness of proposed modules: Due to the differences in model architecture between the current semi-supervised change detection methods referred to and ours, we did not rush to verify the superiority of our method at first. Instead, we conducted model architecture experiments first, using the classical Mean-Teacher architecture. In addition to setting hyperparameters for it to control the weight of unsupervised losses, the rest of the data augmentations and CD network remained the same as [27] and [25]. The gain of this semi-supervised framework compared with the single model and the two-branch network with shared weights is very obvious, and it can almost approach the previous optimal performance. This also demonstrates the validity of the principle of perturbed consistency and the parameter integration, on the basis of which we explore the adaptive training mechanism. Therefore, we separately integrated our proposed AdaEMA and AdaFusion into the MT framework and achieved average improvement of 1.2 and 5.1 on $IoU^{c}$ respectively, as shown in Table IV. * in the table indicates that in adapt, an adaptive judgment operation is performed to determine whether the fusion is performed, and a random selection strategy is used when selecting the fusion region, and a huge gap between the two is evident. Finally, the two adaptive modules contain the complete method, and better results are obtained on the basis of individual modules, which shows that the two modules proposed by us are decoupled, and the model architecture is reasonable.
TABLE IV: Ablation study of our proposed AdaSemiCD on LEVIR-CD Dataset. AEMA and AF denotes our AdaEMA, AdaFusion module respectively. And AF* represents the fusion region are randomly selected.
Method 5% 10% 20% 40%
$IoU^{c}$ OA $IoU^{c}$ OA $IoU^{c}$ OA $IoU^{c}$ OA
Sup. only 61.0 97.60 66.8 98.13 72.3 98.44 74.9 98.60
MT-EMA 67.1 (+6.1) 98.14 (+0.54) 75.0 (+8.2) 98.63 (+0.50) 76.6 (+4.3) 98.71 (+0.27) 77.0 (+2.1) 98.73 (+0.13)
MT-AEMA 68.9 (+7.9) 98.23 (+0.63) 76.1 (+9.3) 98.66 (+0.53) 77.7 (+5.4) 98.78 (+0.34) 77.8 (+2.9) 98.78 (+0.18)
(MT-EMA)+AF* 72.0 (+11.0) 98.43 (+0.83) 76.8 (+10.0) 98.72 (+0.59) 77.5 (+5.2) 98.74 (+0.30) 78.5 (+3.6) 98.80 (+0.20)
(MT-EMA)+AF 77.0 (+16.0) 98.72 (+1.12) 78.8 (+12.0) 98.83 (+0.70) 80.4 (+8.1) 98.91 (+0.47) 80.0 (+5.1) 98.90 (+0.30)
AdaSemiCD 77.7 (+16.7) 98.78 (+1.18) 79.4 (+12.6) 98.87 (+0.74) 80.3 (+8.0) 98.92 (+0.48) 80.6 (+5.7) 98.93 (+0.33)
Hyperparameters in Ramp-up: The Ramp-up process has a significant influence on the performance of our AdaSemiCD on SSCD. Therefore, we conduct experiments on the selection of two hyperparameters ($\gamma$ and $w_{max}$) a and b that control the Ramp-up process. As shown in Table V, our method achieves the best performance on the three datasets under the combination of parameters (0.1, 10), (0.1, 0.1), and (0.1, 1.0) respectively. Moreover, our method is sensitive to this hyperparameter, and inappropriate parameter selection will cause large performance attenuation. This is because our method conducts supervised training on labeled samples and unsupervised training on unlabeled samples at the same time. If the relationship between the two cannot be properly balanced, overfitting on labeled samples or excessive noise interference from unlabeled samples will be caused. All of our remaining experiments were performed at this hyperparameter setting, and the hyperparameters of the compared methods were consistent with the best choices in their original paper.
[…]
$\gamma$ $w_{max}$ LEVIR-CD WHU-CD CDD
$IoU^{c}$ OA $IoU^{c}$ OA $IoU^{c}$ OA
0 0 (Sup.Only) 66.8 98.13 55.7 97.53 67.9 95.46
0.05 1.0 67.2 98.17 53.8 97.02 74.4 96.35
0.1 1.0 71.8 98.32 61.0 98.10 77.3 96.89
0.3 1.0 69.9 98.26 59.4 98.03 76.2 96.56
0.5 1.0 68.7 98.15 60.9 98.25 76.3 96.58
1.0 1.0 67.3 98.13 60.5 98.18 72.3 95.98
0.1 0.1 65.2 97.60 70.8 98.70 71.6 95.77
0.1 0.5 68.3 98.14 66.9 98.54 75.8 96.67
0.1 5.0 73.9 98.75 60.1 98.00 69.1 95.51
0.1 10.0 79.4 98.87 52.4 97.40 68.2 95.50
0.1 30.0 71.9 98.42 50.34 97.12 65.4 95.20
IV-D Complexity Analysis
Since we utilize the same architecture with FPA and RCR, the number of training parameters (46.85M) and computational amount (585.85 GFLOPs) were the same. The variations in parameters, FLOPs, and the time required for training and inference are presented in Table VI. Our AdaSemiCD is comparable to other methods with reference count, FLOPs, and inference time. The main reason for the longer training time was that it took about 0.3s for each iteration to generate and evaluate pseudo-labels twice, while it only took about 0.006s and 0.03s for fusion and EMA parameters updating, respectively. Moreover, our model balances performance and time consumption, offering a notable performance benefit with only a slight increase in time.
[…]
Method Params(M) FLOPs(G) Training Time(s) Inference Time(ms) $IoU^{c}$
Sup.Only   46.85  585.85    77     56 61.0
AdvEnt[68]   46.85  585.85    405     63 66.1
s4GAN[37]   46.85  585.85    585     58 64.0
SemiCDNet[26]   46.85  585.85    408     75 67.6
RCR[25]   46.85  585.85    742     59 72.5
FPA[27]   46.85  585.85    727     68 73.7
AdaSemiCD   46.85  585.85    915     67 77.7
Oracle   46.85  585.85    293     55 77.9
IV-E Limitation Analysis and Future Work
Although our approach is effective in most scenarios, we also found in our experiments that there are certain limitations in some cases. As shown in Fig. 5, our approach struggled and even performed worse than the supervised approach when faced with the detection of areas of leptosomatic variation, such as highways, rivers, and so on, especially from the example in the upper right corner, it can be seen that our method has great advantages in the identification and detection of blocky change areas. This is because when our AdaFusion transforms samples, it is the fusion of patch to patch, which easily truncates this changing region that throughout the entire image, and there is too little real data with annotations, resulting in no effective knowledge to be learned.
Therefore, in future research work, more targeted fusion methods could designed to avoid such information loss. For example, a changing region is regarded as an instance and sample fusion is carried out at the instance level. In this way, the characteristics of different change types can be preserved to a great extent while noise is removed, additional supervisory information is introduced and sample diversity is expanded. We will also actively explore the application of this adaptive mechanism to other semi-supervised tasks and frameworks.
Figure 5: Examples of common error cases in the CDD dataset.
V Conclusion
In this study, we present AdaSemiCD, a flexible semi-supervised framework for change detection. This framework assesses the quality of pseudo-labels on unlabeled training samples and implements adaptive modifications based on the assessment outcomes, which include sample fusion (AdaFusion), and parameter updates (AdaEMA). Despite the complexity of the scenes, our model successfully identifies the areas of interest with minimal interference during training. Empirical evidence from several publicly available CD datasets attests to the efficacy of our methodology. Looking ahead, this adaptive processing technique holds promise for potential application in other semi-supervised tasks.

### v2

@@
-AdaSemiCD: An Adaptive Semi-Supervised Change Detection Method Based on Pseudo-Label Evaluation
+AdaSemiCD: An Adaptive Semi-supervised Change Detection Method Based on Pseudo Label Evaluation
 Abstract
-Change Detection (CD) is an essential field in remote sensing, with a primary focus on identifying areas of change in bi-temporal image pairs captured at varying intervals of the same region by a satellite.
-The data annotation process for the CD task is both time-consuming and labor-intensive.
-To make better use of the scarce labeled data and abundant unlabeled data, we present an adaptive dynamic semi-supervised learning method, AdaSemiCD, to improve the use of pseudo-labels and optimize the training process.
+Change detection(CD) is an essential field in remote sensing, with a primary focus on identifying areas of change in bitemporal image pairs captured at varying intervals of the same region.
+The data annotation process for CD tasks is both time-consuming and labor-intensive.
+To better utilize the scarce labeled data and abundant unlabeled data, we introduce an adaptive semi-supervised learning method, AdaSemiCD, to improve pseudo-label usage and optimize the training process.
 Initially, due to the extreme class imbalance inherent in CD, the model is more inclined to focus on the background class, and it is easy to confuse the boundary of the target object.
-Considering these two points, we develop a measurable evaluation metric for pseudo-labels that enhances the representation of information entropy by class rebalancing and amplification of confusing areas to give a larger weight to prospects change objects.
-Subsequently, to enhance the reliability of sample-wise pseudo-labels, we introduce the AdaFusion module, which is capable of dynamically identifying the most uncertain region and substituting it with more trustworthy content.
+Considering these two points, we develop a measurable evaluation metric for pseudo-labels that enhances the representation of information entropy by class rebalancing and amplification of ambiguous areas, assigning greater weights to prospective change objects.
+Subsequently, to enhance the reliability of samplewise pseudo-labels, we introduce the AdaFusion module, to dynamically identifying the most uncertain region and substituting it with more trustworthy content.
 Lastly, to ensure better training stability, we introduce the AdaEMA module, which updates the teacher model using only batches of trusted samples.
-Experimental results from LEVIR-CD, WHU-CD, and CDD datasets validate the efficacy and universality of our proposed adaptive training framework.
+Experimental results on ten public CD datasets validate the efficacy and generalizability of our proposed adaptive training framework.
 Index Terms:
-Change Detection, Semi-supervised Learning, Remote Sensing, Mean-Teacher
+Pseudo label, Semi-supervised Learning, Change Detection, Mean Teacher, Adaptive Learning
 publicationid: pubid: $0$–$00$ 2021 IEEE
 I Introduction
-Change detection (CD) has emerged as a significant research focus within the field of remote sensing in recent years.
-Its objective is to identify regions of interest that have experienced alterations in bi-temporal image pairs captured at varying times of the same geographical area using satellite technology.
-This method plays a crucial role in remote sensing data analysis and is of particular importance in various civilian sectors such as urban planning [1, 2], rural land management [3, 4], and disaster assessment [5, 6].
-Figure 1: Comparison between SOTA semi-supervised methods and our AdaSemiCD on the CDD Dataset.
-Given that the process of accurately annotating masks for change detection tasks is notably labor intensive, direct application of traditional supervised learning approaches, such as CNN [7, 3, 8, 9, 4] and Transformers [10, 11, 12], to a limited set of labeled data often results in limited performance.
-In response to these challenges, researchers have investigated a range of approaches such as Self-supervised learning(SSL)[13, 14],Unsupervised Change Detection (USCD) [15, 16, 17], Weakly Supervised Change Detection (WSCD) [18, 19, 20], and sample generation strategies [21, 22, 23, 24].
-Although WSCD is cost-efficient, it is reliant on incomplete or inaccurate labels, which can introduce substantial errors and unpredictably noisy data.
-USCD, on the other hand, does not require labeled data and utilizes the intrinsic patterns present in the data, but it often faces challenges when tackling specific tasks like classification or detection.
-Sample generation strategies, which include data augmentation [23], generative adversarial networks (GANs) [22], and diffusion models [24], frequently necessitate the simulation or synthesis of additional data.
+Change detection (CD) has emerged as a significant research focus within the field of remote sensing.
+Its objective is to identify regions of interest that have experienced alterations in bi-temporal image pairs captured at varying times of the same geographical area.
+This method plays a crucial role in remote sensing data analysis and is particularly important in various civilian sectors, such as urban planning [1, 2], rural land management [3, 4], and disaster assessment [5, 6].
+Given that the process of accurately annotating masks for change detection tasks is notably labor-intensive, direct application of traditional supervised learning approaches, such as convolutional neural networks (CNN) [7, 8] and Transformers [9, 10], to a limited set of labeled data often results in limited performance.
+In response to these challenges, researchers have explored a range of approaches such as weakly supervised change detection (WSCD) [11, 12], unsupervised change detection (USCD) [13, 14], and semi-supervised change detection (SSCD) [15, 16].
+Although WSCD is cost-efficient, it relies on incomplete or inaccurate labels, which can introduce significant errors and unpredictable noise.
+USCD, on the other hand, does not require labeled data and leverages the intrinsic patterns present in the data; however, it often faces challenges when tackling specific tasks like classification or detection.
+Some methods would adopt sample generation strategies [17, 18, 19, 20], which include data augmentation [19], generative adversarial networks (GAN) [18], and diffusion models [20], frequently necessitate the simulation or synthesis of additional data.
 However, when dealing with limited available samples, these methods may encounter constraints due to insufficient diversity in the generated data, which can diminish the model’s ability to generalize.
-As a result, semi-supervised change detection (SSCD) [25, 26, 27] emerges as a potentially more effective solution.
-The paradigm of semi-supervised learning [28, 29, 30] is to enhance CD performance by leveraging the limited available labels and the large volume of unlabeled samples.
-Usually, researchers generate pseudo-labels for the unlabeled data to act as guidance during training.
+As a result, SSCD [21, 22, 23] emerges as a potentially more effective solution.
+The paradigm of semi-supervised learning (SSL) [24, 25, 26] aims to enhance CD performance by leveraging the limited available labeled data and the large volume of unlabeled samples.
+Typically, researchers generate pseudo-labels for the unlabeled data to act as guidance during training.
 These pseudo-labels are often temporary predictions with higher probabilities.
-The most prevalent approach is the Mean-Teacher [31] framework, which employ a teacher model to generate pseudo-labels that serve as guidance for the student model during the training process.
+The most widely used approach is the mean-teacher (MT) [27] framework, which employs a teacher model to generate pseudo-labels that serve as guidance for the student model during the training process.
 The teacher model is subsequently updated using the exponential moving average (EMA) [28] of the student model.
-By training on a mix of limited data with actual labels and abundant data with pseudo-labels, the student model can learn more significant features, leading to noticeable enhancements in performance.
+The student model benefits from training on a mix of limited labeled data along with ample pseudo-labeled data, enabling it to identify more important features and resulting in marked enhancements in performance.
 While these methods produce acceptable outcomes, significant problems persist: the model indiscriminately treats all samples, irrespective of their quality, and the training process lacks flexibility.
-Firstly, it’s evident that unlabeled samples may not always function as efficient ‘teachers’.
-Models frequently encounter difficulties in generating reliable high-quality pseudo-labels for intricate samples, which in turn introduces extra noise that can misguide the model’s training.
+Firstly, it is evident that unlabeled samples may not always function as efficient ‘teachers’.
+Models frequently encounter difficulties in generating reliable high-quality pseudo-labels for intricate samples, which in turn introduces extra noise that can mislead the model’s training.
 Subsequently, the EMA updating process does not take into account the quality of samples.
-Given that training batches can be biased or contain noise, dynamically determining the training update timing could contribute to the stability of the training process.
+Given that training batches may be biased or contain noise [29], dynamically determining the timing of training updates could contribute to the stability of the training process.
 These factors underscore the need for a more precise supervisory approach, failing which it could negatively impact the model’s training.
-In this study, we introduce an adaptive dynamic learning strategy, AdaSemiCD, designed to improve the accuracy of pseudo-labels and streamline the training process.
-Our framework incorporates the traditional semi-supervised training approach, augmented by two innovative functional modules, AdaFusion and AdaEMA.
-Initially, we utilize AdaFusion to suppress noise at the individual sample level, thereby enhancing the accuracy of pseudo-labels.
-Contrary to previous methods like Augseg [33] or CutMix [34] that relied on entirely random fusion regions, our AdaFusion technique proactively identifies the most uncertain region and substitutes it with reliable content from either labeled datasets or unlabeled datasets of superior quality.
+In this study, we introduce an adaptive learning strategy, AdaSemiCD, designed to improve the accuracy of pseudo-labels and streamline the training process.
+Our framework builds upon the traditional semi-supervised training approach, augmented by two innovative functional modules, AdaFusion and AdaEMA.
+Initially, AdaFusion is employed to suppress noise at the individual sample level, thereby enhancing the accuracy of pseudo-labels.
+Contrary to previous methods like Augseg [30] or CutMix [31] that relied on entirely random fusion regions, our AdaFusion technique proactively identifies the most uncertain region and substitutes them with reliable content from either labeled datasets or unlabeled datasets of superior quality.
 Following this, we dynamically adjust the rate of parameter updates in the teacher-student model via AdaEMA to ensure improved stability.
-Although the traditional EMA effectively mitigates fluctuations in model parameters, thereby boosting stability, it persists in uniformly updating after each training iteration, overlooking the model’s varying learning outcomes across different iterations when handling a range of training samples.
-If unlabeled samples contain an abundance of erroneous information, it can misdirect the model’s training.
-Therefore, our AdaEMA introduces an adaptive selection process for model-level parameter updates, enabling the model to fully integrate superior parameters.
-In summary, the main contributions of this paper are as follows.
+Although the traditional EMA effectively mitigates fluctuations in model parameters, thereby boosting stability, it persists in uniformly updating after each training iteration, neglecting the model’s varying learning outcomes across different iterations when handling a range of training samples.
+If unlabeled samples contain excessive erroneous information, it can misdirect the model’s training.
+Therefore, our AdaEMA introduces an adaptive selection process for model-level parameter updates, allowing the model to fully integrate superior parameters.
+The main contributions of this paper are as follows:
   • We propose an adaptive SSCD framework named AdaSemiCD, which dynamically improves the pseudo-labels as well as adjusts the training procedure with pseudo label quality assessment.
@@
   • We propose an AdaEMA parameter update strategy, which updates the teacher model with a batch-wise pseudo-labels improving assessment.
-  • Experimental results on publicly available datasets LEVIR-CD, WHU-CD, and CDD demonstrate the effectiveness of our method.
+  • Experimental results on ten publicly available datasets demonstrate the effectiveness of our method.
 II Related Work
 II-A Semi-supervised Learning
-The technique of semi-supervised learning (SSL) applies supervised learning to a small subset of labeled data and unsupervised learning to a larger pool of unlabeled data.
-SSL is mainly categorized into three strategies: consistent regularization (CR), self-training, and holistic methods.
-The last strategy combines the first two strategies within an SSL framework.
+Semi-supervised learning involves applying supervised learning on a limited amount of labeled data while employing unsupervised learning on a vast set of unlabeled data.
+SSL is typically divided into three strategies: consistent regularization (CR), self-training, and holistic methods, with the latter integrating the first two strategies within an SSL framework.
 CR techniques are grounded in the concept of perturbed consistency, which utilizes the coherence between the model’s output after varying degrees of perturbed input data as a training constraint.
-The three consistent regularization frameworks, in ascending order, including: the ${\Pi}$-model [35], the Temporal-ensembling model [35], and the Mean-teacher (MT) model [32].
+The three typical consistent regularization frameworks consist of the ${\Pi}$-model [32], the Temporal-ensembling model [32], and the mean-teacher model [28].
 The ${\Pi}$-model’s double-branch network shares the weight; the Temporal-ensembling model amalgamates all the outputs in the time series, with each image’s pseudo-labels being the EMA of the previously generated results; the MT model carries out this smoothing operation at the model parameters level.
-This model has found application in subsequent semi-supervised research across various domains, such as Active-Teacher for semi-supervised object detection [36], [37, 38, 33, 39] for semi-supervised general semantic segmentation,  [40] for image classification, and  [41][42] for semi-supervised medical image segmentation.
-There has also been explored in the area of perturbation design, with [43] and [25] examining the image-level perturbation and feature-level perturbation of CR respectively.
+This model has found application in subsequent semi-supervised research across various domains, such as Active-Teacher for semi-supervised object detection [33], [34, 35, 36] for semi-supervised general semantic segmentation, [37] for image classification, and [38, 39] for semi-supervised medical image segmentation.
+There has also been explored in the area of perturbation design, with [40] and [21] examining the image-level and feature-level perturbation of CR respectively.
 In the realm of self-training methods, the authenticity of pseudo-labels is of paramount importance.
 This has led to extensive research into the effective selection of high-quality pseudo-labels for supervised learning.
+Feng et al.[41] proposed a method that incrementally adds labeled instances, reduces class bias via Synthetic minority over-sampling technique, and adaptively selects the optimal number of instances to enhance classifier performance.
 [42] employs a set probability threshold as a selection standard.
@@
 [44] uses a constant entropy value as the filtering limit.
-More often than not, CR and self-training are not used independently, but rather in conjunction, forming a comprehensive approach to semi-supervised learning.
-This is exemplified in [38, 33, 39, 40] and [45], as partially discussed in earlier sections.
+CR and self-training are commonly employed together rather than in isolation, creating a holistic strategy for semi-supervised learning.
+This is illustrated in [34, 30, 36, 37] and [43], as partially described in previous sections.
 II-B Semi-supervised Change Detection
 Since annotating a large number of images for CD is time-consuming, recent methods mainly focus on the SSCD.
-In the realm of CR techniques, the incorporation of the mean teacher model in CD was first introduced by Bousias et al.
+In the realm of CR techniques, the incorporation of the mean-teacher model in CD was first introduced by Bousias et al.
 [45].
@@
 [48] propose a dual-task SSCD framework that combines building segmentation and change detection, two closely related downstream tasks.
-They devised a novel consistency constraint between the two change detection masks produced by the Siamese segmentation network and the change detection network.
+They devised a novel consistency constraint between the two change detection masks produced by the siamese segmentation network and the CD network.
 Bandara et al.
 [21] explored feature-based perturbations of the regularization term, applying various data perturbations at the feature level to expand the distribution space of consistency constraints.
-This approach fully leverages the information embedded in unlabeled samples, and in recent work, Zhang et al.
+This approach fully leverages the information embedded in unlabeled samples.
+In recent work, Zhang et al.
 [23] imposed two constraints of class consistency and feature consistency on unlabeled datasets.
 By aligning the feature representations of unlabeled samples on varying and invariant classes, the model could learn from a feature space that is closer to the real distribution, this contributed to the renewal of the best performance record at that time.
-Other methods predominantly utilize Generative Adversarial Network (GAN), which were initially proposed by Ian J.
-Goodfellow in 2014[51].
-Some methods use GAN to learn the feature distribution space close to the real labeled data[26, 52, 53, 54]; The other part uses GAN to generate data samples[55, 56]; The more noteworthy research in the recent work is [18] proposed a new paradigm for change detection, unifying unsupervised, weakly supervised, regionally supervised, and fully supervised change detection into an end-to-end framework.
-While these endevavors have reported great success in SSCD, the highly unstable training of GAN makes hyperparameter adjustment become s challenging.
-Additionally, the issue of gradient disapperance frequently arises during the training phase.
-Morever, the strong discriminatory capaality of the discriminator can lead to imbalanced performance between the generator and discriminator of the GAN if no additional training techniques are implementrd.
-Consquently, achieving the idealized optimal scenario is challenging, making the pritical application of this method somewhat intricate.
-So we continue the semi-supervised framework of consistent and self-teinging.
+Some approaches primarily employ generative adversarial networks to produce data samples and understand feature distributions that approximate real labeled data [22, 49, 50].
+Although these efforts have shown notable achievements in SSCD, the notoriously erratic nature of GAN training presents difficulties with hyperparameter tuning.
+Furthermore, the occurrence of gradient vanishing is a common challenge during training.
+Moreover, the discriminator’s robust ability to differentiate can cause an imbalance between the generator and discriminator’s performance within the GAN unless additional training strategies are applied.
+As a result, reaching an ideal balance is demanding, complicating the practical application of this method.
+Therefore, we continue the semi-supervised framework leveraging consistency and self-training techniques.
 III METHODOLOGY
 Fig.
-2 provides a comprehensive summary of our AdaSemiCD framework, aiming to enhance the SSCD performance by utilizing the scarce labels and vast quantity of unlabeled samples.
-We commence with a broad introduction of the framework, followed by an in-depth explanation of the uncertainty map used to assess our pseudo-labels, and finally present the specifics of AdaFusion and AdaEMA.
-Figure 2: The overview of the proposed AdaSemiCD framework, and details of two novel modules.
-Our approach is based on the common teacher-student training pipeline for SSCD, and we suggest utilizing the AdaFusion module to produce samples with increased reliability and the AdaEMA module to enhance the efficiency of EMA update times.
+1 provides a comprehensive summary of our AdaSemiCD framework, aiming to enhance the SSCD performance by leveraging scarce labeled data and a vast quantity of unlabeled samples.
+We commence with a general introduction of the framework, followed by an detailed explanation of the uncertainty map used to assess our pseudo-labels, and finally present the specifics of AdaFusion and AdaEMA.
+Figure 1: The overview of the proposed AdaSemiCD framework and details of two adaptive modules.
+Our approach is based on the common MT training pipeline for SSCD, and we suggest utilizing the AdaFusion module to produce samples with increased reliability and the AdaEMA module to enhance the efficiency of EMA update.
 III-A Overview of the AdaSemiCD Framework
-The task of semi-supervised change detection can be described as follows: Given a labeled dataset ${D}_{l}=\left\{\left\{\smash{x_{a,i}^{l},x_{b,i}^{l}}\right\},y_{i}^{l}\right\}_{i=1}^{m}$ and an unlabeled dataset ${D}_{u}=\left\{x_{a,j}^{u},x_{b,j}^{u}\right\}_{j=1}^{n}$, where $\left\{\left\{\smash{x_{a,i}^{l},x_{b,i}^{l}}\right\},y_{i}^{l}\right\}$ represents the i-th pair of labeled image pairs and their ground truth, and $\left\{\smash{x_{a,i}^{u},x_{b,i}^{u}}\right\}$ represents the i-th pair of unlabeled image pairs, subscripts a and b are used to identify the images before and after phases respectively, and the number of samples with and without labeled datasets is n and m, $n\gg m$, the model $M$ is expected to not only extract meaningful information from ${D}_{l}$ but also capture a wider range of features by utilizing numerous unlabeled training samples in ${D}_{u}$ to improve the model’s generalization capabilities.
-Model architecture: In this study, we employ the widely-used teacher-student framework for SSCD tasks.
-The network is composed of two components, a student model $M_{stu}$ and a teacher model $M_{tea}$ that shares the same structure.
+The process of semi-supervised change detection is broadly outlined as follows: We are given a labeled dataset, denoted as ${D}_{l}=\left\{\left\{\smash{x_{a,i}^{l},x_{b,i}^{l}}\right\},y_{i}^{l}\right\}_{i=1}^{m}$, and an unlabeled dataset ${D}_{u}=\left\{x_{a,j}^{u},x_{b,j}^{u}\right\}_{j=1}^{n}$.
+Here, $\left\{\left\{\smash{x_{a,i}^{l},x_{b,i}^{l}}\right\},y_{i}^{l}\right\}$ illustrates the $i$-th pair of labeled images alongside their corresponding true labels, while $\left\{\smash{x_{a,j}^{u},x_{b,j}^{u}}\right\}$ refers to the $j$-th pair of unlabeled images.
+The subscript $a$ signifies the image from the ‘Pre’-event period, and $b$ signifies the ‘Post’-event period.
+Importantly, the quantity of samples with labels and those without are $m$ and $n$, with $n$ considerably larger than $m$.
+The aim is for model $M$ to not only derive key insights from ${D}_{l}$ but also to enhance its feature extraction capability using the extensive collection of unlabeled samples in ${D}_{u}$, thus boosting the model’s generalization potential.
+Ordinarily, samples are subject to either weak augmentation $A_{w}(\cdot)$ or strong augmentation $A_{s}(\cdot)$ prior to being passed into the network to guarantee superior generalization performance.
+Model architecture: In this study, we employ the widely used MT framework for SSCD tasks.
+Two integral parts make up the network: a student model, denoted as $M_{stu}$, and a teacher model, represented as $M_{tea}$.
+Both components possess an identical architecture.
 The student model is trained to extract significant features from a small number of labeled samples and a large volume of unlabeled samples, with the aid of optimization via gradient descent methods.
-Conversely, the teacher model $M_{tea}$ generates pseudo-labels to guide the student in assimilating unlabeled data, and the teacher model is refreshed using EMA methods.
-Usually, samples undergo either weak $A_{w}(\cdot)$ or strong $A_{s}(\cdot)$ augmentation processes prior to being input into the network to maintain high-quality generalization performance.
-Objectives: The objective is to minimize the supervised loss $\mathcal{L}_{s}$ on $D_{l}$ while also ensuring consistency on disturbed $D_{u}$ with a minimal $\mathcal{L}_{u}$.
+Conversely, the teacher model $M_{tea}$ generates pseudo-labels to guide the student in assimilating unlabeled data, and it is updated using the EMA method.
+Objectives: The objective is to minimize the supervised loss $\mathcal{L}_{s}$ on $D_{l}$ while ensuring consistency on the disturbed $D_{u}$ with a minimal $\mathcal{L}_{u}$.
 During training, samples are fed into the network in randomly shuffled batches, $\mathcal{B}_{l}$ and $\mathcal{B}_{u}$.
-The loss $\mathcal{L}_{s}$ for labeled samples within a batch $B_{l}$ is calculated as the cross entropy (CE) of ground truth $y_{i}^{l}$ and it’s prediction $p_{i}^{l}$:
+The loss $\mathcal{L}_{s}$ for labeled samples within a batch $\mathcal{B}_{l}$ is calculated as the cross entropy(CE) of ground truth $y_{i}^{l}$ and its prediction $p_{i}^{l}$:
 $\mathcal{L}_{s}=\frac{1}{\left|\mathcal{B}_{l}\right|}\sum_{i=1}^{\left|\mathcal{B}_{l}\right|}\mathrm{CE}\left({p}_{i}^{l},y_{i}^{l}\right),$ (1)
-where $|\mathcal{B}_{l}|$ represents the mini-batch’s size, ${p}_{i}^{l}$ represents the CD result of the $i$-th pair images.
+where $|\mathcal{B}_{l}|$ represents the mini-batch size, ${p}_{i}^{l}$ represents the change detection result for the $i$-th pair of images.
 The loss $\mathcal{L}_{u}$ for unlabeled samples within a batch $B_{u}$ is quite similar.
-Here we use the pseudo-labels from $M_{tea}$ as supervision for the prediction from $M_{stu}$.
-And $\mathcal{L}_{u}$ is calculated as:
+Here, we use the pseudo-labels from $M_{tea}$ to supervise the predictions from $M_{stu}$.
+The loss $\mathcal{L}_{u}$ is calculated as follows:
 $\mathcal{L}_{u}=\frac{1}{\left|\mathcal{B}_{u}\right|}\sum_{j=1}^{\left|\mathcal{B}_{u}\right|}\mathrm{CE}\left(p_{s,j}^{u},{p}_{w,j}^{u}\right),$ (2)
-where $|\mathcal{B}_{u}|$ represents the size of unlabeled image mini-batch, ${p}_{w,j}^{u}=M_{tea}^{\theta^{\prime}}(w_{a,j}^{u},w_{b,j}^{u})$ represents the detection result of the teacher model on the change of the i-th pair of unlabeled images weakly augmented, and ${p}_{s,j}^{u}={M_{stu}^{\theta}(s_{a,j}^{u},s_{b,j}^{u})}$ represents the detection result of the teacher model on the change of the i-th pair of unlabeled images by further strongly augmented.
-To sum up, the total loss for the training procedure of AdaSemiCD is:
-$\mathcal{L}=\mathcal{L}_{s}+\lambda(\cdot)\mathcal{L}_{u}$ (3)
-Previous studies[40, 41] have employed a constant value for $\lambda(\cdot)$.
-However, we argue that this approach is not suitable for single-stage semi-supervised learning algorithms.
-During the initial training phases, the pseudo-labels generated by our model for unlabeled samples are highly unreliable; excessive reliance on unsupervised training at this stage can introduce significant noise.
-Conversely, as training progresses into the middle and later stages, the model increasingly learns from the limited labeled data, leading to an improvement in the quality of the pseudo-labels.
-This justifies a reduction in the proportion of supervised training relative to unsupervised training.
-To mitigate overfitting and enhance the feature space, it is essential to implement a controlled process to dynamically adjust the contribution of these two components in the loss function.
-As the pseudo-labels in the starting stage are quite unstable, we use a ramp-up process to control the training pace of the unlabeled part.
-$\lambda(\cdot)$ denotes the time varing weight of unsupervised loss, which is a function changing with training step and used to dynamically adjust the proportion of supervised and unsupervised training in different stages.
+where $|\mathcal{B}_{u}|$ represents the size of the unlabeled image mini-batch.
+${p}_{w,j}^{u}=M_{tea}^{\theta}(w_{a,j}^{u},w_{b,j}^{u})$ and ${p}_{s,j}^{u}=M_{stu}^{\theta}(s_{a,j}^{u},s_{b,j}^{u})$ denote the change detection outcomes from the teacher model for the $j$-th pair of weak and strong augmentations of unlabeled images, respectively.
+To summarize, the overall loss associated with the AdaSemiCD training process is defined as:
+$\mathcal{L}=\mathcal{L}_{s}+\lambda(\cdot)\mathcal{L}_{u}.$ (3)
+$\lambda$ is the weight of unsupervised loss, which is typically set to a constant value [37, 38].
+For SSCD, we contend that using a constant $\lambda$ could potentially disrupt the training procedure.
+In the early stages of training, the pseudo-labels for unlabeled data tend to be highly unreliable, and excessive dependence on unsupervised training at this point can inject substantial noise.
+Conversely, as training progresses into the intermediate and final phases, the model enhances its ability to generate high-quality pseudo labels, diminishing the importance of the limited labeled dataset.
+This shift warrants a gradual decrease in the emphasis on supervised training compared to unsupervised training over time.
+To prevent overfitting and refine the feature space, it is crucial to have a systematic approach that dynamically modifies the balance between these two elements in the loss function.
+We employ a ramp-up approach, defining $\lambda(\cdot)$ as a function that adapts through the course of training, thereby dynamically modulating the balance between supervised and unsupervised training throughout various phases.
 $\lambda(\cdot)=w_{max}\times e^{-\phi\times\left(1-iter_{cur}/iter_{max}\right)^{2}}$ (4)
-where $w_{max}$ and $\phi$ are hyperparameters, $w_{max}$ representing the maximum weight of unsupervised loss, $\phi$ is used to control the severity of the ramp-up; $iter_{cur}$ represents the current iteration cycles; $iter_{max}$ is the total number of ramp-up cycles, is obtained by multiplying $\gamma$ and the total number of training iterations, and $0\textless\gamma\textless 1.0$; After the ramp-up process, the weight of unsupervised losses remains $w_{max}$ and does not change.
-In the early stage of training, this weight is relatively low, and unsupervised training plays a negligible role, but in the middle and later stages of training, this weight gradually increases and the unsupervised loss after weighting exceeds the supervised loss, so unsupervised training accounts for the main part.
-Training strategy: The student network’s parameters $\theta$ are adjusted by Stochastic Gradient Descent(SGD) technique to reduce the overall loss $\mathcal{L}$, while the teacher network’s parameters $\theta^{\prime}$ are updated by the exponential moving average of the student model’s parameters $\theta$ over the time series, as depicted in Eq.
-5.
-The hyperparameter $\beta$ acts as a momentum parameter.
-The larger the value of $\beta$, the broader the moving metric average window becomes.
-Typically, $\beta$ is chosen to be a number near 1.0, for instance, 0.996 in this study.
-$\theta^{\prime}=\beta\theta^{\prime}+(1-\beta)\theta$ (5)
-Proposed modules: The essence of semi-supervised learning lies in the quality of pseudo-labels.
-Nevertheless, it’s apparent that the previously mentioned procedure does not take into account the diverse effects of individual samples on training.
-This paper concentrates on two elements that are directly related to the generation of pseudo-labels: the initial pair of unlabeled images, and the efficiency of the pseudo-label-generation network, also known as the teacher model, in identifying changes.
-To offer more dependable supervisory data to unlabeled information and reduce training uncertainty, we’ve designed an adaptive training strategy to tackle these two key issues.
-Our initial proposal is to develop a metric that can measure the uncertainty of pseudo-labels, serving as the basis for adaptive modifications.
-Following that, we suggest making adaptive changes at the image level to the unlabeled training samples and suitably integrating reliable contents.
-Furthermore, it includes applying adaptive and selective EMA updates to the teacher network during the training phase to limit variations, thus producing more consistent and superior quality pseudo-labels.
-These efforts are elaborated in the subsequent sections.
+$iter_{max}=\gamma\times iter_{t}otal$ (5)
+Here, $w_{max}$ represents the maximum weight value of the unsupervised loss, and $\phi$ controls the severity of the ramp-up.
+$iter_{cur}$ represents the current iteration cycles; $iter_{max}$ is the total number of ramp-up cycles, calculated by multiplying $\gamma$ (where $10<\gamma<$ 1) by the total number of training iterations, as shown in 5.
+After the ramp-up process, the weight of the unsupervised loss stabilizes at $w_{max}$ and no longer changes.
+In the early stages of training, this weight is relatively low, so unsupervised training plays a negligible role, but in the middle and later stages of training, this weight gradually increases and the unsupervised loss after weighting exceeds the supervised loss, making unsupervised training the dominant factor.
+Training strategy: To reduce the total loss $\mathcal{L}$, the parameters of the student network, denoted as $\theta_{stu}$, are refined via Stochastic Gradient Descent (SGD).
+Concurrently, the teacher network updates its parameters $\theta_{tea}$ using an exponential moving average calculated from the student network’s parameters $\theta_{stu}$ over a time sequence:
+$\theta_{tea}=\beta\theta_{tea}+(1-\beta)\theta_{stu}$ (6)
+The hyperparameter $\beta$ acts as a momentum factor, where a higher $\beta$ value leads to a broader moving average window.
+Generally, $\beta$ is selected to be near 1.0; in this study, for instance, $\beta$ is set at 0.996.
+Proposed modules: The effectiveness of semi-supervised learning depends largely on the quality of pseudo-labels.
+Nevertheless, it is clear that the previously mentioned procedure does not take into account the varying impact of individual samples on training.
+This paper concentrates on two critical aspects that are directly related to the generation of pseudo-labels: the initial pair of unlabeled images, and the efficiency of the pseudo-label-generation network, aka the teacher model, in identifying changes.
+To offer more reliable supervisory signals to unlabeled information and reduce training uncertainty, we propose an adaptive training strategy to tackle these two key issues.
+Our initial proposal is to develop a metric that quantifies the uncertainty of pseudo-labels, serving as the basis for adaptive modifications.
+Following that, we recommend applying adaptive modifications at the image level to the unlabeled training samples and suitably integrating reliable contents.
+Additionally, we apply adaptive and selective EMA updates to the teacher network during the training phase to minimize variations, ensuring more consistent and higher-quality pseudo-labels.
+The details of these methods are discussed in the following sections.
 III-B Pseudo-label Qualification Metric
 To enhance the effectiveness of pseudo-labels by accurately gauging their quality, a crucial step is the implementation of an evaluation metric.
-This metric will be instrumental in identifying reliable labels and determining the utility of each training sample.
-Unlike the scenario with labeled image pairs where real labels serve as a benchmark for assessing the model’s inference quality through average intersection over union, F1 scores, and so forth, pseudo-labels lack any such reference points and can only be compared to themselves.
-Therefore, we have devised a quantifiable evaluation metric specifically for pseudo-labels, designed to increase the total information entropy while taking into account factors like category imbalance and areas of confusion.
-To evaluate the quality of pseudo-labels, a frequently employed technique for this is the computation of information entropy [57],
+This metric plays a key role in identifying reliable labels and determining the value of each training sample.
+Unlike labeled image pairs, where true labels serve as a benchmark for assessing the model’s performance, pseudo-labels lack such reference points and can only be compared to themselves.
+To evaluate the quality of pseudo-labels, a common technique is the computation of information entropy [51],
 $E(x_{i})=-P\left(x_{i}\right)\log_{2}P\left(x_{i}\right),$ (7)
 where $P\left(x_{i}\right)$ is the output probability of a trained model on sample $x_{i}$.
-Our conviction is that a lower information entropy in the predicted value corresponds to a greater trustworthiness of the prediction outcome.
-On the contrary, a higher information entropy leads to a more varied prediction outcome, a more evenly distributed prediction probability on the same pixel, and a diminished distinction in the category.
+We propose that reduced information entropy in the predicted values implies greater reliability of the prediction.
+In contrast, increased information entropy points to a prediction with more uncertainty, indicating a more balanced probability distribution across the pixels and resulting in decreased confidence in the model’s predictions.
 Figure 2: Class statistics of common CD datasets. The proportion of the changed/unchanged categories is extremely unbalanced.
-In our change detection task, the direct application of entropy often does not yield good results due to the significant challenge posed by category imbalance.
-As illustrated in Fig.
-3, ten publicly available change detection datasets were evaluated, namely HRCUS-CD[58], LEVIR-CD[59], LEVIR-CD+[59],WHU-CD[60], EGY-CD[61], GZ-CD[26], CL-CD[62], CDD[63], SYSU-CD[64] and DSIFN-CD[65].
-The ratios of the changed and unchanged categories are highly skewed.
-This significant category imbalance can lead to our model inadequately learning the target categories during training while disproportionately capturing feature distributions from background categories.
-As a result, there is an increased tendency to classify pixels as belonging to the background class during inference.
-Consequently, although the overall accuracy (OA) of the model appears high, the performance on the Intersection over Union (IoU) across various categories remains unsatisfactory.
-To minimize the influence of this category imbalance when assessing the quality of pseudo-labels, we assigned different weights to the two categories when computing information entropy as Eq.
-7.
+In the CD task, directly applying entropy may not yield the optimal results due to the significant challenge posed by class imbalance, as illustrated in Fig.
+2.
+The ratios of changed to unchanged categories are highly skewed.
+This imbalance may lead our model to learn the target categories inadequately during training while disproportionately capturing feature distributions from the background categories.
+As a result, the model tends to classify pixels as background during inference.
+To reduce the influence of this class imbalance when assessing the quality of pseudo-labels, we assign different weights to the two categories when computing information entropy,
 $E^{\prime}\left(X\right)=w_{1}\times E\left(x_{i}\right)[0]+w_{0}\times E\left(x_{i}\right)[1],$ (8)
-where $w_{0}$ and $w_{1}$ represent the proportions of pixels in the current batch that belong to the unchanged and changed categories, respectively.
-The formulas for their calculation are given by Eq.
-8 and Eq.
-9.
+where $w_{0}$ and $w_{1}$ represent the proportions of pixels in the current batch that belong to the unchanged and changed categories, respectively. The formulas for their calculation are:
 $w_{0}=\frac{\sum_{i=1}^{|Bu|}\sum_{k=1}^{H\times W}{(P(x_{i})==0)}}{{|Bu|}\times H\times W}$ (9)
 $w_{1}=\frac{\sum_{i=1}^{|Bu|}\sum_{k=1}^{H\times W}{(P(x_{i})==1)}}{{|Bu|}\times H\times W}$ (10)
-Moreover, pixels from diverse areas contribute in varying manners.
-Pixels of greater importance, like those in boundary, target, and regions similar to the background, who are offenly predicted with large uncertainty, should receive more emphasis.
-Initially, we enhance the impact of these essential regions by computing the absolute value of the disparity between the prediction probabilities of the two categories.
+Furthermore, pixels from different regions play distinct roles.
+Pixels that are more crucial, like those found on edges, targets, or areas resembling the background—frequently predicted with higher uncertainty—should be prioritized.
+To achieve this, we start by enhancing the significance of these key regions by calculating the absolute difference between the prediction probabilities of the two classes,
 $D\left(x_{i}\right)=abs(P\left(x_{i}\right)[1]-P\left(x_{i}\right)[0])$ (11)
-Here, $abs$ signifies absolute value, a measure taken to prevent the inconsistencies that may arise from varying changes both prior to and following a phase.
-Then doing pixel-level product operation with information entropy, we get the uncertainty map of image $(x_{i})$,
+In this context, $abs$ represents the absolute value function, employed to prevent inconsistencies that could occur due to differing changes before and after a phase.
+By carrying out a pixel-wise multiplication with the information entropy, we derive the uncertainty map for the image $x_{i}$,
 $U\left(x_{i}\right)=1-D\left(x_{i}\right)\cdot E^{\prime}\left(x_{i}\right)$ (12)
-Clearly, in pixel locations with low information entropy, this process won’t result in notable alterations, whereas in pixel locations with high information entropy, the absolute value of the predictive probability post-subtraction will be significantly reduced, and the immediate amalgamation will diminish its importance.
-Consequently, we adopt the inverse of rectified information entropy as its depiction.
-To sum up, with the goal of evaluating the quality of pseudo-labels for change detection, we put forward a measurable calculation metric $U$ to gauge uncertainty, which enhances the overall information entropy considering factors like category imbalance and focus on areas of confusion.
+In areas characterized by low information entropy, this procedure is unlikely to cause significant changes. However, in regions where information entropy is high, the operation enhances the effect.
+To summarize, aiming to assess the quality of pseudo-labels in change detection, we propose a quantifiable metric $U$ to evaluate uncertainty.
+This metric improves overall information entropy by considering elements such as class imbalance and regions of confusion.
 III-C AdaFusion: Adaptive Sample Fusion
-Image fusion is frequently employed to augment samples and boost generalization, with CutMix[34] and MixUp[66] being typical examples.
-In this study, our aim is to leverage image fusion techniques to exclude the unreliable areas in a training sample.
-This procedure involves two steps: region selection, which decides the location of the operation, and image selection, which decides the content to be used.
+Image fusion is widely used to augment samples and enhance generalization, with CutMix[31] and MixUp[52] being typical examples.
+In this study, our aim is to apply image fusion techniques to exclude unreliable areas from the training samples.
+This process consists of two steps: region selection, which determines the location for the operation, and image selection, which specifies the content to be used.
 Adaptive selection of fusion region.
-Contrary to the conventional CutMix technique, which arbitrarily chooses blending regions, our approach is more refined.
-Initially, we would select a bounding box with random size, then slide the window, and choose the area where the total of uncertainty is highest as the region to be combined, using Eq.
-11.
-These regions are frequently boundaries or intricate areas where it’s challenging to identify targets.
-The CutMix fusion can guarantee the variety of samples and simultaneously decrease the unsupervised noise.
+Contrary to the conventional CutMix technique, which arbitrarily selects blending regions, our approach is more refined.
+First, we initialize a bounding box of random size, then slide the window to identify the area with the highest total uncertainty using Eq.
+(12), as the region to be fused.
+These regions typically include boundaries or complex areas where target identification is difficult.
+The selection approach guarantees a varied sample and, at the same time, decreases unsupervised noise.
 Adaptive selection of fusion contents.
-For the region of maximum uncertainty, we have the option to choose a substitute from its corresponding samples from the label section $B_{l}$ or the unlabeled section with higher probabilities $B_{u}$.
-This strategy prevents overuse of only labeled samples and further mitigates the risk of overfitting.
-The focus of the content here is confined to the contrast between labeled and unlabeled image pairs, rather than a more detailed distinction (such as the level of the image or the level of the object).
-This is because we suggest that too much human intervention could potentially undermine the model’s capacity to generalize.
-The choice of fusion content is mainly accomplished by adaptively adjusting a threshold that determines whether to fuse with a labeled image pair.
-We directly use the computed uncertainty as the threshold, and if the randomly generated probability exceeds the total uncertainty of the sample, the labeled images in the training batch are randomly selected as the fusion content.
-If not, other unlabeled image pairs in the batch are randomly selected.
-It is easy to understand that the higher the sample’s reliability, the better the quality of the pseudo-label.
-We consider it reliable enough to avoid the necessity for fusion with the limited labeled images, which could potentially increase the likelihood of overfitting.
-Samples with higher uncertainty come with more noise, and the addition of new content can significantly reduce the noise density.
+For the region of maximum uncertainty, we can select a substitute from either the corresponding samples in the labeled set $B_{l}$ or the unlabeled set $B_{u}$ with higher reliability.
+This strategy helps prevent overreliance on labeled samples and further mitigates the risk of overfitting.
+The selection of fusion content is guided by an adaptive threshold that determines whether to fuse with a labeled image pair.
+We directly use the computed uncertainty as the threshold, and if the randomly generated probability exceeds the total uncertainty of the sample, labeled images in the training batch are selected as the fusion content.
+Otherwise, other unlabeled image pairs of higher quality are chosen.
+It is clear that the higher the reliability of a sample, the better the quality of the pseudo-label.
+If a sample is considered reliable enough, we avoid fusing it with limited labeled images, which could increase the risk of overfitting.
+Samples with higher uncertainty contain more noise, and incorporating new content is expected to reduce the noise effectively.
 III-D AdaEMA: Adaptive EMA Update Strategy
-Within the teacher-student framework, the teacher model is obtained by integrating the moving exponential average of the student model across the time series.
-This approach is more stable and reliable compared to dual models with shared weights.
-However, our expectation for Mean-Teacher is to reach the optimal state of coevolution, where the teacher model is a cumulative representation of the evolving student model.
-The crucial element in achieving this co-evolution is whether the student model progresses or regresses each time the teacher model gets updated.
-So, how can we assess the state of the student model?
-This question is inherently related to the validation phase of the training process.
-Typically, after training for several epochs, we evaluate the current model, feed samples from the validation set into the model for inference, and compare the outcomes with the actual labels to determine accuracy.
-However, if we perform such validation after every iteration, it becomes highly computationally intensive, leading to a significant increase in training time.
-Lowering the number of samples in the validation set (to just a few pairs) might address the feasibility issue, but it also presents new challenges when the sample size is too small to thoroughly assess the model’s performance.
-So, is there a middle ground that can merge these two concepts?
+Within the MT framework, the teacher model’s update process involves incorporating the student’s exponential moving average over time.
+Our goal is to reach a state of optimal coevolution, where the teacher acts as an aggregate reflection of the student model’s progress.
+A pivotal factor in this coevolution is whether the student model advances or declines with each update of the teacher.
+Evaluating the student model’s state is closely tied to the training process’s validation phase.
+Typically, after several training epochs, we evaluate the model by using samples from the validation set, performing inference, and comparing the output against true labels to determine accuracy.
+Conducting this validation after every iteration incurs a significant computational cost and extends training duration.
+One may address this by reducing the validation set size to just a few pairs; however, this risks insufficiently assessing model performance if the sample is too small.
+The challenge lies in striking a balance between these considerations.
+During each training phase, we commence by updating the student model, denoted as $M_{stu}^{\theta}$, according to the strategy outlined in section III-A, which results in the updated model $M_{stu}^{\theta^{\prime}}$.
+Subsequently, we assess both the modified student model, $M_{stu}^{\theta^{\prime}}$, and the teacher model, $M_{tea}^{\theta}$ on the current set of unlabeled training samples, $\mathcal{B}_{u}$.
+Using Eq.
+(12), we calculate the corresponding uncertainty maps, referred to as $U_{tea}$ and $U_{stu}$.
+The changes in the student model’s development are captured through fluctuations in uncertainty,
+$\varepsilon=\frac{\sum{U_{stu}}-\sum{U_{tea}}}{|B_{u}|}.$ (13)
+Additionally, we introduce a probability $\tau$ to regulate model updates according to the reliability of the student model, defined as
+$\tau=\left\{\begin{array}[]{cl}\frac{1}{iter^{2}+\epsilon}&,\quad\varepsilon\leq 0\\ 1.0&,\quad\varepsilon>0\\ \end{array}\right.,$ (14)
+Here, $iter$ denotes the current iteration count and $\epsilon=1e-5$.
+When $\varepsilon$ is greater than zero, the student model progresses, allowing a straightforward update of $M_{tea}^{\theta}$.
+In contrast, if $\varepsilon$ is zero or less, the student model encounters either regression or fluctuation.
+To gauge the probability of altering $M_{tea}^{\theta}$, we incorporate some randomness.
+The adaptive updating details are provided in Algorithm 1.
 Algorithm 1 The AdaEMA algorithm.
-0:    Student model $M_{stu}^{\theta}$, Teacher model $M_{tea}^{\theta^{\prime}}$ The set of training samples for current batch =$\mathcal{B}=\left\{\mathcal{B}_{l},\mathcal{B}_{u}\right\}$;
-0:    Updated Teacher model, $M_{tea}^{\hat{\theta}}$;
-1:  Calculate the loss $\mathcal{L}_{s}$ on labeled samples $\mathcal{B}_{l}$ with Eq. 1
-2:  Calculate the loss $\mathcal{L}_{u}$ on unlabeled samples $\mathcal{B}_{u}$ with Eq. 2
-3:  Update the student model $M_{stu}^{\theta}$ to $M_{stu}^{\theta^{\prime}}$ by SGD to minimize the total loss following Eq. 3
-4:  Calculate the uncertainty ${U}_{tea}$ of pseudo-labels generated on $\mathcal{B}_{u}$ by the teacher model $M_{tea}^{\theta^{\prime}}$according to Eq.11
-5:  Calculate the uncertainty ${U}_{stu}$ on $\mathcal{B}_{u}$ by updated student model $M_{stu}^{\theta^{\prime}}$ according to Eq.11
-6:  Calculate the upper bound on the probability of update $\tau$ according to Eq. 12 and Eq. 13;
+0:    Student model $M_{stu}^{\theta}$, Teacher model $M_{tea}^{\theta}$ The set of training samples for the current batch, =$\mathcal{B}=\left\{\mathcal{B}_{l},\mathcal{B}_{u}\right\}$
+0:    Updated Teacher model, $M_{tea}^{\theta^{\prime}}$;
+1:  Calculate the supervised loss $\mathcal{L}_{s}$ on labeled samples $\mathcal{B}_{l}$ using Eq. (1);
+2:  Calculate the unsupervised loss $\mathcal{L}_{u}$ on unlabeled samples $\mathcal{B}_{u}$ using Eq. (2);
+3:  Update the student model $M_{stu}^{\theta}$ to $M_{stu}^{\theta^{\prime}}$ using SGD to minimize the total loss, as described in Eq. (3).
+4:  Calculate the uncertainty ${U}_{tea}$ of the pseudo-labels generated on $\mathcal{B}_{u}$ by the teacher model $M_{tea}^{\theta}$ as described in Eq. (12);
+5:  Calculate the uncertainty ${U}_{stu}$ of the pseudo-labels generated on $\mathcal{B}_{u}$ by updated student model $M_{stu}^{\theta^{\prime}}$ as described in Eq. (12);
+6:  Calculate the upper bound of the update probability $\tau$ according to Eq. (13) and Eq. (14);
 7:  if $random(\cdot)<\tau$ then
-8:     Update the teacher model to $M_{tea}^{\hat{\theta}}$ by Eq. 5
+8:     Update the teacher model to $M_{tea}^{\theta^{\prime}}$ by EMA;
 9:  else
-10:     $M_{tea}^{\hat{\theta}}=M_{tea}^{\theta^{\prime}}$
+10:     $M_{tea}^{\theta^{\prime}}=M_{tea}^{\theta}$
 11:  end if
-12:  return  $M_{tea}^{\hat{\theta}}$.
-At each training phase, we initially update the student model’s parameters, denoted as $M_{stu}^{\theta}$, following the training strategy described in section III-A, resulting in $M_{stu}^{\theta^{\prime}}$.
-Next, we evaluate both the updated student model $M_{stu}^{\theta^{\prime}}$ and the teacher model $M_{tea}^{\theta^{\prime}}$ on the current batch of unlabeled training samples $\mathcal{B}_{u}$.
-We determine the ensuing uncertainty map using Eq.
-11, and denote these as $U_{tea}$ and $U_{stu}$.
-The progression is marked by changes in uncertainty, as illustrated in Eq.
-12.
-$\varepsilon=\frac{\sum{U_{stu}}-\sum{U_{tea}}}{|B_{u}|}.$ (12)
-If $\varepsilon$ is below zero, the model either degrades or oscillates.
-In contrast, when $\varepsilon$ is zero or higher, the model progresses.
-Ultimately, only the student models that have evolved successfully are selected to refresh the teacher model, with the maximum probability of update $\tau$ being derived as presented in Eq.
-13.
-$\tau=\left\{\begin{array}[]{cl}\frac{1}{iter^{2}+\epsilon}&,\quad\varepsilon\leq 0\\ 1.0&,\quad\varepsilon>0\\ \end{array}\right.,$ (13)
-Here, $\epsilon=1e-5$ is used to prevent the divisor from being zero, and $iter$ denotes the present iteration count.
-If the model parameters are not selected for updating the Teacher model, they are not discarded outright.
-Instead, they are given a lower probability of being updated.
-With this upper bound regulation (ranging from 0 to 1), some randomness is introduced to determine the likelihood of updating the teacher network’s parameters.
-Further details are available in Algorithm 1.
-IV Experiement
-This section pertains to the experimental validation we performed.
-Initially, we present the assessment standards for experimental outcomes and the experimental datasets.
-Subsequently, we executed ablation experiments to validate the significance of each component in the methodology.
-Ultimately, we draw comparisons with the cutting-edge semi-supervised change detection techniques, using measurable experimental metrics and perceptible visual outcomes.
+12:  return  $M_{tea}^{\theta^{\prime}}$.
+IV Experiment
 IV-A Experimental Setup
 IV-A1 Datasets
-Our method is empirically tested on ten benchmark datasets, namely the LEVIR-CD [59], LEVIR-CD+ [59], WHU-CD[60], EGY-CD[61], HRCUS-CD[58], Change Detection Dataset(CDD)[63], GZ-CD[26], DSIFN-CD[65], SYSU-CD[64] and CL-CD[62].
-Among them, these datasets cover different resolutions (0.03m-2.0m), different data sizes (2400-20000 pairs), different annotation categories (binary or multiclass), and different time spans between image pairs (1-16 years), as summarized in the Table I.
+Figure 3: Typical examples of ten change detection datasets we use, with Pre-event image, Post-event image, and Ground Truth.
+Our method is empirically tested on ten benchmark datasets, namely the LEVIR-CD [53], LEVIR-CD+ [53], WHU-CD[54], EGY-CD[55], HRCUS-CD[56], CDD [57], GZ-CD[22], DSIFN-CD[58], SYSU-CD[59], and CL-CD[60].
+As summarized in Table I, these datasets cover different resolutions (0.03m-2.0m), different data sizes (1-20000 pairs), different annotation categories (binary or multiclass), and different time spans (2-16 years).
+Fig.
+3 shows some classic sample images for each dataset.
 TABLE I: Ten publicly available change detection datasets used in the experiment.
@@
 EGY-BCD [55] 0.25m 256 $\times$ 256 6091 2015 to 2022 Link
-Building HRCUS-CD [58] 0.5m 256 $\times$ 256 11388 Varing Link
+Binary HRCUS-CD [56] 0.5m 256 $\times$ 256 11388 Varing Link
 CDD[57] 0.03m-1.0m 256 $\times$ 256 16000 Varing Link
@@
 IV-A2 Evaluation Metrics
-To facilitate a straightforward comparison with the latest cutting-edge techniques, we employed intersection over union (IoU) scores and overall accuracy (OA) on evaluation datasets.
-Given the severe imbalance in CD categories and our primary concentration on the changed area, we use the IoU of the change category $IoU^{c}$, the calculation formula are as follows:
-$\mathrm{IoU}=\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FP}+\mathrm{FN}}$ (14)
-$\mathrm{OA}=\frac{\mathrm{TP}+\mathrm{TN}}{\mathrm{TP}+\mathrm{FP}+\mathrm{TN}+\mathrm{FN}}$ (15)
+For easy comparison with the most advanced techniques, we utilized overall accuracy (OA) to assess general performance.
+Due to the significant imbalance in CD categories and our main focus on the altered area, we applied the intersection over union for the changed category $IoU^{c}$.
+The calculation formulas are provided below:
+$IoU^{c}=TP/(TP+FP+FN)$ (15)
+$OA=(TP+TN)/(TP+FP+FN+TN)$ (16)
 Where $\mathrm{TP}$ represents the positive sample correctly predicted (the correct changing pixel), $\mathrm{TN}$ refers to the negative sample correctly predicted (the correct unchanged pixel), and $\mathrm{FP}$ denotes the positive sample wrongly predicted (the unchanged pixel wrongly detected), $\mathrm{FN}$ represents the negative sample wrongly predicted (the pixel missed as the unchanged pixel).
 For both metrics, the larger the value, the better the change detection performance of the model.
-In the change detection task, OA is high overall due to binary classification and unbalanced category envy.
 IV-A3 Implementation Details
-The common components of this experiment were tested under the same conditions as the methods we intend to compare them with.
-To prevent network structure from causing interference, we utilized ResNet50+PPM as our change detection network follow [25] and [27].
-We set the initial learning rate at 0.01, which linearly decreases to 1e-4 with a momentum of 0.9, and trained all methods using the SGD optimizer.
-All models underwent 80 epochs of training, with each batch having a mini-batch size of 8 for both labeled and unlabeled training sets.
-Furthermore, all the enhancements employed were identical to those used in FPA[27], which include random flip, random resize ranging from 0.5 to 2.0, and random crop among the weak augmentations, along with the nine strong augmentations cited in [67].
-Given the limited amount of labeled data, the model is highly sensitive to the pseudo-labels’ filtering threshold, leading us to set 0.95 as the threshold for all models in the experiments conducted on the ten datasets and we follow the experience in [25] to set $\phi$ to 5.
-All our experiments were carried out using PyTorch on four NVIDIA GeForce RTX 3090 GPUs.
+We employ ResNet50+PPM as our change detection framework, as referenced in [21] and [23].
+The learning rate starts at 0.01 and decreases linearly to 1e-4, with momentum kept at 0.9.
+Training of all competing approaches is conducted using the SGD optimizer over 80 epochs.
+For both labeled and unlabeled data, the mini-batch size is set at 8.
+Moreover, the augmentations applied are the same as those in FPA [23], which consist of weak augmentations such as random flip, random resizing from 0.5 to 2.0, and random cropping, along with nine robust augmentations [61].
+A pseudo label threshold of 0.95 is adopted for all models in our studies.
+When calculating loss, $\phi$ is fixed at 5.
+The entire set of experiments was executed using PyTorch on four NVIDIA GeForce RTX 3090 GPUs.
 IV-B Compare with the SOTA Methods
-In order to prove the superiority of our proposed method, we compare it with several of the most advanced semi-supervised change detection methods.
-[26], [25] and [27] refresh the last semi-supervised CD performance in chronological order.
-In addition, we also compare with two semi-supervised semantic segmentation modding methods[37] and [68].
+To demonstrate the advantages of our proposed method, we have conducted a comparison with a number of leading semi-supervised change detection techniques [22, 21, 23, 35, 62].
 IV-B1 Quantitative Results
-We conducted a complete experiment on ten datasets, Table II and Table III report our experimental results on the Building CD datasets and mutilclass CD datasets respectively.
-Sup.Only is the result of supervised training only on a limited subset of labeled dataset, while Oracle is the result of fully supervised training on the entire training dataset.
-It is worth noting that our method achieves the state-of-the-art performance in almost all partition settings across the all datasets.
-To be sure, all the semi-supervised CD methods performed better than the supervised methods in the corresponding partition setting in most cases, which proves the effectiveness of the semi-supervised methods in learning from a large number of unlabeled training samples, at the same time, the lead achieved by our method proves that our adaptive design for how to learn from unlabeled samples to more correct guidance is of practical significance.
+Tables II and III display our experimental outcomes for both the Building CD datasets and the multi-class CD datasets.
+The ‘Sup.
+only’ results denote supervised training outcomes from a restricted portion of the labeled dataset, whereas ‘Oracle’ represents full supervision results using the entire training dataset.
+It’s notable that our approach achieves SOTA performance in nearly every partition configuration across these datasets.
+Remarkably, in most scenarios, all semi-supervised CD approaches outperformed the supervised ones within the same partition setting, confirming the efficacy of semi-supervised methods in leveraging numerous unlabeled training samples.
+Additionally, the superiority of our method highlights the practical importance of our adaptive strategy for learning more effectively from unlabeled data.
+Building CD Datasets: As presented in Table II, the proposed AdaSemiCD framework demonstrates outstanding performance across nearly all building CD datasets.
+Notably, on LEVIR-CD, LEVIR-CD+, and WHU-CD datasets, AdaSemiCD achieves a significant improvement in $IoU^{c}$, with gains of 3.1, 1.3, and 1.7 percentage points, respectively.
+On EGY-CD and HRCUS-CD datasets, AdaSemiCD maintains an overall leading position with average improvements of 0.75 and 0.4 points, respectively.
+However, performance on certain configurations is marginally lower than SOTA methods.
+This limitation can be attributed to the unique challenges posed by these datasets: the EGY-CD dataset suffers from overexposure, which often blends buildings with the background, increasing the difficulty of accurate discrimination.
+Similarly, the HRCUS-CD dataset is affected by vegetation occluding the changing areas of some buildings.
+These challenges hinder the model’s ability to generate reliable pseudo-labels when trained with a limited number of real labels.
 […]
-In addition, from the experimental results, the overall accuracy(OA) of all methods on these binary change detection datasets is very high, because the background class accounts for the majority.
-OA is often directly corresponding to $IoU_{c}$, and a slight improvement in OA is reflected in a relatively significant improvement in $IoU_{c}$.
-Our AdaSemiCD improves OA in almost all datasets because it greatly eliminates noise and the guidance of false signals.
+Furthermore, as indicated by the experimental results, the overall accuracy of all methods on these binary change detection datasets is notably high, primarily due to the dominance of the background class.
+OA often correlates closely with $IoU^{c}$, where even a slight improvement in OA translates into a relatively significant improvement in $IoU^{c}$.
+Our AdaSemiCD consistently enhances OA across nearly all datasets, as it effectively reduces noise and minimizes the influence of false signals, leading to more accurate predictions.
 TABLE II: The average quantitative metrics of different CD methods on building change detection datasets.
@@
 SemiCDNet[22] 52.4 96.00 57.9 96.31 62.8 96.95 63.8 97.19
-RCR[25] 58.1 96.50 61.9 96.77 63.9 97.08 64.2 97.18
+RCR[21] 58.1 96.50 59.9 96.77 63.9 97.08 64.2 97.18
 FPA[23] 57.5 96.52 60.1 96.86 65.2 97.25 65.7 97.34
@@
 AdaSemiCD 36.9 80.46 39.2 82.94 41.1 85.45 45.1 87.12
-Oracle $IoU^{c}$=37.1 and OA=86.82
+Oracle $IoU^{c}$=58.1 and OA=90.82
 SYSU-CD Sup. only 62.9 89.57 64.4 90.18 66.0 90.82 66.4 90.93
@@
 AdaSemiCD 67.5 91.16 68.7 91.59 70.1 92.03 69.9 91.90
-Oracle $IoU^{c}$=77.9 and OA=98.77
+Oracle $IoU^{c}$=68.2 and OA=91.64
 CL-CD Sup. only 18.1 91.90 31.4 92.42 37.2 93.32 45.9 94.98
@@
 AdaSemiCD 30.6 92.52 33.5 92.40 41.6 94.21 49.1 95.85
-Oracle $IoU^{c}$=77.9 and OA=98.77
-Multiclass CD Datasets: As shown in TableIII, our AdaSemiCD achieved SOTA performance on almost all datasets and all settings, just the improvement was less than on the binary building change detection datasets.
-This is due to the complexity of the task itself, and one of the most obvious is that in some cases, some of the previous semi-supervised change detection methods failed and performed worse than the baseline under full supervision.
-However, our semi-supervised CD approach can still leverage additional unlabeled samples to improve performance.
-Another point worth noting is that in these multiclass change detection datasets, the overall accuracy of all methods is reduced, and OA and $IoU_{c}$ are no longer directly corresponding.
-One of the phenomena is that $IoU_{c}$ reaches the highest, but OA does not, which is due to its different emphasis.
-OA pays more attention to the global accuracy, that is, the detection of background categories.
-$IoU_{c}$ is more concerned with the detection of the target class, which is more important in CD than the background class.
-As you can see in DSIFN-CD, our approach has achieved good results on the $IoU_{c}$.
+Oracle $IoU^{c}$=50.1 and OA=95.66
+Multiclass CD Datasets: As shown in Table III, our proposed AdaSemiCD demonstrates optimal performance in the majority of scenarios, despite the increased complexity of multi-category change detection tasks compared to binary building change detection tasks.
+Notably, optimal results were achieved across all experimental settings on the CDD dataset, with $IoU^{c}$ and OA increasing by an average of 1.8 percentage points and 0.25 percentage points respectively, which can be attributed to its exceptionally accurate annotations, as illustrated in Fig.
+3.
+This high-quality labeling is particularly advantageous for pseudo-label-based semi-supervised methods.
+Furthermore, AdaSemiCD exhibits overall performance improvements on the SYSU-CD and CL-CD datasets.
+However, its performance is slightly below the SOTA results in certain experimental configurations.
+This can be attributed to the inherent challenges in identifying some mountain and land changes present in these datasets.
+These changes are often large-scale, and misclassification in such regions may result in fluctuations in performance metrics.
+An additional significant observation is that, when the overall accuracy declines, the correlation between OA and $IoU^{c}$ becomes less straightforward.
+In particular, there are instances where $IoU^{c}$ reaches its maximum value, while OA does not, underscoring their different focuses.
+OA is mainly concerned with overall accuracy, highlighting the detection of background categories, whereas $IoU^{c}$ prioritizes the precise detection of target classes, which is more vital for change detection tasks than identifying background categories.
+This difference is apparent in the DSIFN-CD dataset, where our method demonstrates strong performance on the $IoU^{c}$ metric.
+[…]
 IV-B2 Qualitative Results
-To provide a clearer illustration of the change detection capability of our proposed method, we conducted visualizations on three different datasets.
 Fig.
-4 showcases an example of the visualizations on the WHU-CD test sets.
-It is apparent that on both datasets, our approach has notably mitigated the common issues of missed and false detections.
+4 and Fig.
+5 showcase some examples of the visualizations on the test sets of building and mutilclass CD datasets respectively, in which the area selected in the box is the error-prone area.
+It is apparent that on those datasets, our approach has notably mitigated the common issues of missed and false detections.
 In challenging scenarios, our method could still effectively identify the areas of change that were of interest to us.
-Regarding the image pairs in the lower part of Fig.
-4, there is significant interference information present between the images.
-Many of the detected changes are unrelated to the buildings of interest.
+Some failure cases of the detected changes are unrelated to the buildings of interest.
 The absence of adequate supervision information makes it challenging to mitigate such interference, leading to decreased model performance.
 Additionally, the task is further complicated by the detection of small and densely changing areas, which proves to be difficult for the model.
-Alternative semi-supervised models may have been trained to detect errors because of guidance issues in this complicated scenario.
-Consequently, it can be seen that in such instances the semi-supervised method performs even worse than the supervised method trained solely on labeled data.
-Our model, on the other hand, adaptively excludes these noises during the training phase and integrates parameters from the model that has completed its evolution through the training process.
-In these intricate cases, the quality of pseudo-labels improves progressively until achieving a reliable level, furnishing an accurate signal for model training.
-Therefore, it is evident that our model excels in these complex regions (highlighted by boxes in Fig.
-4).
-This marks the foundation of our proposed method.
-Figure 4: Visualizations of different models on the WHU-CD dataset. Pre and Post represent the image of the front and back temporal respectively, and GT is the Ground Truth.
+[…]
+[…]
+In complex scenarios, alternative semi-supervised models may encounter guidance issues, leading them to inadvertently amplify errors during training.
+As a result, semi-supervised methods may perform worse than supervised methods trained exclusively on labeled data in such situations.
+Our model, however, adaptively mitigates these noises during the training phase by dynamically excluding them and integrating parameters from a progressively refined model throughout the training process.
+In these intricate cases, the quality of pseudo-labels is incrementally improved until they reach a reliable standard, thereby providing an accurate signal for model training.
+Consequently, our model demonstrates superior performance in these challenging regions, as highlighted by the boxes in Fig.
+4 and Fig.
+5.
+This adaptive approach forms the cornerstone of our proposed method.
 IV-C Ablation Study
-In this section, we conduct some ablation studies, mainly to verify the role of each module in our proposed approach and to explore some of our hyperparameter selections in the experiment.
 Effectiveness of proposed modules: Due to the differences in model architecture between the current semi-supervised change detection methods referred to and ours, we did not rush to verify the superiority of our method at first.
@@
 AdaSemiCD 77.7 (+16.7) 98.78 (+1.18) 79.4 (+12.6) 98.87 (+0.74) 80.3 (+8.0) 98.92 (+0.48) 80.6 (+5.7) 98.93 (+0.33)
-Hyperparameters in Ramp-up: The Ramp-up process has a significant influence on the performance of our AdaSemiCD on SSCD.
-Therefore, we conduct experiments on the selection of two hyperparameters ($\gamma$ and $w_{max}$) a and b that control the Ramp-up process.
-As shown in Table V, our method achieves the best performance on the three datasets under the combination of parameters (0.1, 10), (0.1, 0.1), and (0.1, 1.0) respectively.
-Moreover, our method is sensitive to this hyperparameter, and inappropriate parameter selection will cause large performance attenuation.
-This is because our method conducts supervised training on labeled samples and unsupervised training on unlabeled samples at the same time.
-If the relationship between the two cannot be properly balanced, overfitting on labeled samples or excessive noise interference from unlabeled samples will be caused.
-All of our remaining experiments were performed at this hyperparameter setting, and the hyperparameters of the compared methods were consistent with the best choices in their original paper.
+Pseudo-label qualification metric: As described in Section III-B, the pseudo-label evaluation metric we proposed is based on information entropy, enhanced by class rebalancing and confusion region amplification.
+To demonstrate the effectiveness, we first evaluated our AdaSemiCD using information entropy as the sole metric, then incrementally incorporated class rebalancing and confusion region amplification, and finally evaluated the full model, which integrates all components.
+The experimental results are presented in Table V.
+When only information entropy is used as the evaluation metric for implementing the adaptive training mechanism, significant performance improvements are achieved, attributed to the AdaFusion and AdaEMA modules we designed.
+Incorporating class rebalancing and confusion region amplification based on information entropy further enhances performance, demonstrating that these improvements enable more accurate identification of pseudo-labels during training, thereby facilitating more precise adaptive operations.
+Notably, the gain from class rebalancing exceeds that from confusion region amplification.
+This is because, with class rebalancing, the evaluation of pseudo labels places greater emphasis on the foreground, somewhat neglecting background identification, thus allowing AdaFusion to focus more on the uncertainty region in the foreground.
+In addition, confusion region amplification proves more beneficial as more labeled data becomes available.
+Finally, using uncertainty, an overall assessment that includes both components, leads to further performance gains, suggesting that these two improvements are independent and can be effectively combined for more accurate assessments.
+TABLE V: Ablation study of pseudo-label qualification metric on LEVIR-CD Dataset.
+Method 5% 10% 20% 40%
+$IoU^{c}$ OA $IoU^{c}$ OA $IoU^{c}$ OA $IoU^{c}$ OA
+Sup. only 61.0 97.60 66.8 98.13 72.3 98.44 74.9 98.60
+Entropy 73.2 (+12.2) 98.50 (+0.90) 77.4 (+10.6) 98.75 (+0.62) 78.6 (+6.3) 98.81 (+0.37) 79.2 (+4.3) 98.89 (+0.29)
+Entropy+rebalance 76.0 (+15.0) 98.63 (+1.03) 78.7 (+11.9) 98.82 (+0.69) 79.6 (+7.3) 98.85 (+0.41) 79.6 (+4.7) 98.86 (+0.26)
+Entropy+confusion 74.6 (+13.6) 98.61 (+1.01) 78.0 (+11.2) 98.79 (+0.66) 79.2 (+6.9) 98.82 (+0.38) 79.8 (+4.9) 98.87 (+0.27)
+Uncertainty 77.7 (+16.7) 98.78 (+1.18) 79.4 (+12.6) 98.87 (+0.74) 80.3 (+8.0) 98.92 (+0.48) 80.6 (+5.7) 98.93 (+0.33)
+We have stored the pseudo-labels generated during the training process, and Fig.
+6 illustrates the $IoU^{c}$ between the pseudo-labels and the corresponding Ground Truth throughout training, both with and without class rebalancing in the pseudo-label evaluation (note: the Ground Truth used here for unlabeled samples are solely for calculating the $IoU^{c}$ with the pseudo-labels and are not involved in any other aspect of the training process).
+It is evident that, although class rebalancing does not directly improve the pseudo-labels, it enables a more accurate evaluation of the pseudo-labels.
+As a result, our adaptive training mechanism enhances the quality of them, bringing them closer to the real labels, which leads to a higher $IoU^{c}$ between them.
+Figure 6: The influence of class rebalancing on the quality of pseudo-labels generated for unlabeled samples during training.
 […]
@@
 $IoU^{c}$ OA $IoU^{c}$ OA $IoU^{c}$ OA
-0 0 (Sup.Only) 66.8 98.13 55.7 97.53 67.9 95.46
+0 0 66.8 98.13 55.7 97.53 67.9 95.46
 0.05 1.0 67.2 98.17 53.8 97.02 74.4 96.35
@@
 Since we utilize the same architecture with FPA and RCR, the number of training parameters (46.85M) and computational amount (585.85 GFLOPs) were the same.
-The variations in parameters, FLOPs, and the time required for training and inference are presented in Table VI.
+The variations in parameters, FLOPs, and the time required for training and inference are presented in Table VII.
 Our AdaSemiCD is comparable to other methods with reference count, FLOPs, and inference time.
@@
 Oracle   46.85  585.85    293     55 77.9
-IV-E Limitation Analysis and Future Work
-Although our approach is effective in most scenarios, we also found in our experiments that there are certain limitations in some cases.
-As shown in Fig.
-5, our approach struggled and even performed worse than the supervised approach when faced with the detection of areas of leptosomatic variation, such as highways, rivers, and so on, especially from the example in the upper right corner, it can be seen that our method has great advantages in the identification and detection of blocky change areas.
-This is because when our AdaFusion transforms samples, it is the fusion of patch to patch, which easily truncates this changing region that throughout the entire image, and there is too little real data with annotations, resulting in no effective knowledge to be learned.
-Therefore, in future research work, more targeted fusion methods could designed to avoid such information loss.
-For example, a changing region is regarded as an instance and sample fusion is carried out at the instance level.
-In this way, the characteristics of different change types can be preserved to a great extent while noise is removed, additional supervisory information is introduced and sample diversity is expanded.
-We will also actively explore the application of this adaptive mechanism to other semi-supervised tasks and frameworks.
-Figure 5: Examples of common error cases in the CDD dataset.
+Hyperparameters in Ramp-up: The Ramp-up process has a significant influence on the performance of our AdaSemiCD on SSCD.
+Therefore, we conduct experiments on the selection of two hyperparameters ($\gamma$ and $w_{max}$) that control the Ramp-up process.
+As shown in Table VI, our method achieves the best performance on the three datasets under the combination of parameters (0.1, 10), (0.1, 0.1), and (0.1, 1.0) respectively.
+Moreover, our method is sensitive to this hyperparameter, and inappropriate parameter selection will cause large performance attenuation.
+This is because our method conducts supervised training on labeled samples and unsupervised training on unlabeled samples at the same time.
+If the relationship between the two cannot be properly balanced, overfitting on labeled samples or excessive noise interference from unlabeled samples will be caused.
+All of our remaining experiments were performed at this hyperparameter setting, and the hyperparameters of the compared methods were consistent with the best choices in their original paper.
 V Conclusion
@@
 Despite the complexity of the scenes, our model successfully identifies the areas of interest with minimal interference during training.
-Empirical evidence from several publicly available CD datasets attests to the efficacy of our methodology.
+Empirical evidence from ten publicly available CD datasets attests to the efficacy of our methodology.
 Looking ahead, this adaptive processing technique holds promise for potential application in other semi-supervised tasks.